### v1

# Implementing Grassroots Logic Programs with Multiagent Transition Systems and AI

Abstract
Grassroots Logic Programs (GLP) is a concurrent logic programming language with variables partitioned into paired readers and writers, conjuring both linear logic and futures/promises: an assignment is produced at most once via the sole occurrence of a writer (promise) and consumed at most once via the sole occurrence of its paired reader (future), and may contain additional readers and/or writers, enabling the concise expression of rich multidirectional communication modalities.
GLP was designed as a language for grassroots platforms—distributed systems with multiple instances that can operate independently of each other and of any global resource, and can coalesce into ever larger instances—with its target architecture being smartphones communicating peer-to-peer. The operational semantics of Concurrent (single-agent) GLP and of multiagent GLP (maGLP) were defined via transition systems/multiagent transition systems, respectively.
Here, we describe the mathematics developed to facilitate the workstation- and smartphone-based implementations of GLP by AI in Dart. We developed dGLP—implementation-ready deterministic operational semantics for single-agent GLP—and proved it correct with respect to the Concurrent GLP operational semantics; dGLP was used by AI as a formal spec, from which it developed a workstation-based implementation of GLP. We developed madGLP—an implementation-ready multiagent operational semantics for maGLP—and proved it correct with respect to the maGLP operational semantics; madGLP is deterministic at the agent level (not at the system level due to communication asynchrony), and is being used by AI as a formal spec from which it develops a smartphone-based implementation of maGLP.
The development process employs three layers of abstraction—the mathematical spec (this paper); an English+code fragments informal spec derived by AI from the math; and the Dart code derived by AI from the informal spec. While the authority is math→spec→Dart, the three layers were harmonized via back-and-forth communication, with programming in Dart revealing gaps and inconsistencies in the informal spec, which in turn revealed gaps and inconsistencies in the math. The latter resulted in several major redesigns of madGLP, the outcome of which is presented here.
keywords:
concurrent logic programming, multiagent systems, transition systems, operational semantics, peer-to-peer, grassroots platforms
1 Introduction
Grassroots Logic Programs (GLP) [shapiro2025glp] is a concurrent logic programming language with variables partitioned into paired readers and writers, conjuring both linear logic [girard1987linear] and futures/promises [friedman1976impact]: an assignment is produced at most once via the sole occurrence of a writer (promise) and consumed at most once via the sole occurrence of its paired reader (future), and may contain additional readers and/or writers, enabling the concise expression of rich multidirectional communication modalities.
GLP was designed as a language for grassroots platforms [shapiro2023grassrootsBA]—distributed systems with multiple instances that can operate independently of each other and of any global resource, and can coalesce into ever larger instances—with its target architecture being smartphones communicating peer-to-peer. The operational semantics of Concurrent (single-agent) GLP and of multiagent GLP (maGLP) were defined via transition systems/multiagent transition systems, respectively [shapiro2021multiagent, shapiro2025glp].
Here, we describe the mathematics developed to facilitate the workstation- and smartphone-based implementations of GLP by AI (Claude) in Dart [dart2024] (a multi-platform smartphone programming language). We developed dGLP—implementation-ready deterministic operational semantics for single-agent GLP—and proved it correct with respect to the Concurrent GLP operational semantics; dGLP was used by AI as a formal spec, from which it developed a workstation-based implementation of GLP. We developed madGLP—an implementation-ready multiagent operational semantics for maGLP—and proved it correct with respect to the maGLP operational semantics; madGLP is deterministic at the agent level (not at the system level due to communication asynchrony), and is being used by AI as a formal spec from which it develops a smartphone-based implementation of maGLP.
The development process employs three layers of abstraction—the mathematical spec (this paper); an English+code fragments informal spec derived by AI from the math; and the Dart code derived by AI from the informal spec. While the authority is math→spec→Dart, the three layers were harmonized via back-and-forth communication, with programming in Dart revealing gaps and inconsistencies in the informal spec, which in turn revealed gaps and inconsistencies in the math. The latter resulted in several major redesigns of madGLP, the outcome of which is presented here. All remarks in the paper were generated by AI, to assist itself in better understanding the math while writing the spec.
A key goal of this paper is to offer a precise interface between the human designer and the AI programmer: formal operational semantics serve as a specification language, constraining the space of legal implementations. The emerging discipline [fowler2025sdd, mundler2025type, blinn2024typed] we advocate and employ is for the human designer and AI to jointly develop and agree upon (1) formal semantics (this paper); (2) informal (English+code) specification (derived by AI); and only then let AI attempt to write (3) Dart code that complies with these formal and informal specs. The current madGLP English+code specification is available as supplementary material.
Paper outline. Section 2 recalls transition systems, implementations, and correctness. Section 3 presents Concurrent GLP: syntax, operational semantics, and dGLP which correctly implements it. Section 4 recalls multiagent transition systems and atomic transactions. Section 5 defines multiagent GLP (maGLP), multiagent deterministic GLP (madGLP), and proves madGLP correctly implements maGLP. Section 6 discusses related work. Section 7 concludes. Appendix A contains single-agent GLP proofs and remarks. Appendix B contains multiagent GLP proofs and remarks. Appendix C provides the complete madGLP specification. Appendix D proves madGLP grassroots. Appendix E provides detailed example traces.
2 Transition Systems and Implementations
This section presents transition systems and implementations among them, providing the mathematical foundation for proving that one transition system correctly implements another.
2.1 Transition Systems
Definition 2.1 (Transition System).
A transition system is a tuple $TS=(C,c_{0},T)$ where $C$ is an arbitrary set of configurations, $c_{0}\in C$ is a designated initial configuration, and $T\subseteq C\times C$ is a transition relation, with transitions written $c\rightarrow c^{\prime}\in T$.
A transition $c\rightarrow c^{\prime}\in T$ is enabled from configuration $c$. A configuration $c$ is terminal if no transitions are enabled from $c$. A computation is a (finite or infinite) sequence of configurations where for each two consecutive configurations $(c,c^{\prime})$ in the sequence, $c\rightarrow c^{\prime}\in T$. We write $c\xrightarrow{*}c^{\prime}$ to denote the existence of a computation (empty if $c=c^{\prime}$) from $c$ to $c^{\prime}$. A run is a computation starting from $c_{0}$, which is complete if it is infinite or ends in a terminal configuration. The outcome of a complete run is determined by a domain-specific function from complete runs to an outcome space.
Liveness. The framework we follow [shapiro2021multiagent] has an elaborate notion of liveness. As GLP is a simple special case, we do not recall it here, but assume the notion of fair runs abstractly and instantiate it concretely when defining GLP (Section 3).
2.2 Implementations
Definition 2.2 (Implementation).
Given two transition systems $TS=(C,c_{0},T)$ (the specification) and $TS^{\prime}=(C^{\prime},c^{\prime}_{0},T^{\prime})$, an implementation of $TS$ by $TS^{\prime}$ is a function $\sigma:C^{\prime}\rightarrow C$ where $\sigma(c^{\prime}_{0})=c_{0}$, in which case the pair $(TS^{\prime},\sigma)$ is referred to as an implementation of $TS$.
Given a computation $r^{\prime}=c^{\prime}_{1}\rightarrow c^{\prime}_{2}\rightarrow\ldots$ of $TS^{\prime}$, $\sigma(r^{\prime})$ is the (possibly empty) computation $\sigma(c^{\prime}_{1})\rightarrow\sigma(c^{\prime}_{2})\rightarrow\ldots$ obtained from the sequence by removing consecutively repetitive elements so that $\sigma(r^{\prime})$ has no stutter transitions of the form $c\rightarrow c$.
Definition 2.3 (Live, Complete, and Correct Implementation).
The implementation $(TS^{\prime},\sigma)$ of $TS$ is:
  • live if $\sigma$ maps every fair $TS^{\prime}$ run $r^{\prime}$ to a fair $TS$ run $\sigma(r^{\prime})$
  • complete if for every complete $TS$ run with outcome $O$, there exists a complete $TS^{\prime}$ run with outcome $O$
  • correct if it is live and complete
3 Concurrent GLP and Its Implementation
3.1 Concurrent GLP
This section presents Concurrent Grassroots Logic Programs (GLP), a concurrent logic programming language in which logic variables are partitioned into paired readers and writers.
GLP Syntax.
Definition 3.1 (GLP Variables).
Let $\mathcal{V}$ denote the set of writers (identifiers beginning with uppercase). Define $\mathcal{V}?=\{X?\mid X\in\mathcal{V}\}$, called readers. The set of all GLP variables is $\hat{\mathcal{V}}=\mathcal{V}\cup\mathcal{V}?$. A writer $X$ and its reader $X?$ form a variable pair.
Definition 3.2 (GLP Terms and Goals).
A term is a variable in $\hat{\mathcal{V}}$, a constant (numbers, strings, or the empty list []), or a compound term $f(T_{1},\ldots,T_{n})$ with functor $f$ and subterms $T_{i}$. Let $\mathcal{T}$ denote the set of all terms. We use standard list notation: [X|Xs] for a list cell, [X1,...,Xn] for finite lists. A term is ground if it contains no variables.
A unit goal is a compound term. A goal is a multiset of unit goals; the empty goal is written true. A clause $A$ :- $B$ has head $A$ (a unit goal) and body $B$ (a goal); a unit clause has empty body.
Definition 3.3 (Single-Occurrence (SO) Invariant).
A term, goal, or clause satisfies the single-occurrence (SO) invariant if every variable occurs in it at most once.
Definition 3.4 (GLP Program).
A clause $C$ satisfies the single-reader/single-writer (SRSW) restriction if it satisfies SO and a variable occurs in $C$ iff its paired variable also occurs in $C$. A GLP program is a finite set of clauses satisfying SRSW; clauses for the same predicate form a procedure. Let $\hat{\mathcal{G}}(P)$ denote the set of goals over $\hat{\mathcal{V}}$ and the vocabulary of $P$.
The SO and SRSW restrictions ensure that GLP eschews unification in favour of simple term matching. The result is linear-logic-like [girard1987linear] and futures/promises-like [friedman1976impact]: each assignment $X:=T$ is produced at most once via the sole occurrence of a writer (promise) $X$, and consumed at most once via the sole occurrence of its paired reader (future) $X?$.
GLP Operational Semantics.
Definition 3.5 (Writers Substitution, Assignment, Readers Substitution and Counterpart).
A GLP writer assignment is a term of the form $X:=T$, $X\in\mathcal{V}$, $T\notin\mathcal{V}$, satisfying SO. Similarly, a GLP reader assignment is a term of the form $X?:=T$, $X?\in\mathcal{V}?$, $T\notin\mathcal{V}$, satisfying SO. A writers (readers) substitution $\sigma$ is the substitution implied by a set of writer (reader) assignments that jointly satisfy SO. Given a writers assignment $X:=T$, its readers counterpart is $X?:=T$, and given a writers substitution $\sigma$, its readers counterpart $\sigma?$ is the readers substitution defined by $X?\sigma?=X\sigma$.
Definition 3.6 (GLP Renaming, Renaming Apart).
A GLP renaming is a substitution $\rho:\hat{\mathcal{V}}\to\hat{\mathcal{V}}$ such that for each $X\in\mathcal{V}$: $X\rho\in\mathcal{V}$ and $X?\rho=(X\rho)?$. Two GLP terms have a variable in common if for some writer $X\in\mathcal{V}$, either $X$ or $X?$ occurs in both. A GLP renaming $\sigma$ renames $T^{\prime}$ apart from $T$ if $T^{\prime}\sigma$ and $T$ have no variable in common.
Definition 3.7 (Writer MGU).
Given two GLP unit goals $A$ and $H$, a writer mgu is a writers substitution $\sigma$ such that $A\sigma=H\sigma$ and $\sigma$ is most general among such substitutions. Unlike standard unification, writer mgu only assigns writers, not readers.
Definition 3.8 (GLP Goal/Clause Reduction).
Given GLP unit goal $A$ and clause $C$, with $H$ :- $B$ being the result of the GLP renaming of $C$ apart from $A$, the GLP reduction of $A$ with $C$ succeeds with result $(B,\sigma)$ if $A$ and $H$ have a writer mgu $\sigma$.
Definition 3.9 (GLP Transition System).
Given a GLP program $P$, an asynchronous resolvent over $P$ is a pair $(G,\sigma)$ where $G\in\hat{\mathcal{G}}(P)$ and $\sigma$ is a readers substitution.
A transition system $GLP(P)=(\mathcal{C},c_{0},\mathcal{T})$ is a GLP transition system over $P$ and initial goal $G_{0}$ satisfying SO if:
  1. 1.
    $\mathcal{C}$ is the set of all asynchronous resolvents over $P$
  2. 2.
    $c_{0}=(G_{0},\emptyset)$
  3. 3.
    $\mathcal{T}$ is the set of all transitions $(G,\sigma)\rightarrow(G^{\prime},\sigma^{\prime})$ satisfying either:
    1. (a)
[…]
    2. (b)
      Communicate: $\{X?:=T\}\in\sigma$, $X?\in G$, $G^{\prime}=G\{X?:=T\}$, and $\sigma^{\prime}=\sigma$
GLP Reduce uses a writer mgu instead of standard unification, and chooses the first applicable clause instead of any clause. The first is the fundamental use of GLP readers for synchronization. The second allows writing fair concurrent programs.
The GLP Communicate rule realizes the use of reader/writer pairs for asynchronous communication: it communicates an assignment from its writer to its paired reader.
GLP has two forms of nondeterminism: the choice of $A\in G$ (and-nondeterminism), and the choice of which Communicate to apply.
Definition 3.10 (Proper Run, Outcome).
A GLP run $\rho:(G_{0},\sigma_{0})\rightarrow\cdots\rightarrow(G_{n},\sigma_{n})$ is proper if for any $1\leq i<n$, a variable that occurs in $G_{i+1}$ but not in $G_{i}$ also does not occur in any $G_{j}$, $j<i$. If proper, the outcome of $\rho$ is $(G_{0}$ :- $G_{n})\sigma_{n}$.
We now discuss GLP persistence and define fair runs, needed to prove a GLP implementation to be live.
The following definition assumes a notion of goal identity—the ability to identify “the same goal” across configurations as its readers are instantiated. It can be formalized by pairing each goal with a unique identifier assigned at spawn time; as this machinery is needed for proving liveness of the implementation of GLP, but not for the actual implementation, we leave it implicit.
Definition 3.11 (Persistent Transition System).
A transition system is persistent if every enabled transition remains enabled until taken.
Lemma 3.12 (GLP Persistence).
GLP is persistent.
See Appendix A for proofs and further remarks.
Definition 3.13 (GLP Fair Run).
A complete GLP run is fair if every transition that becomes enabled is eventually taken.
3.2 Deterministic Implementation of GLP
Motivation. dGLP serves as the formal specification for AI to implement single-agent GLP on workstations. The discipline is: from the mathematical definition (this section), AI derives an English+code specification, and from the spec derives Dart code.
While the GLP transition system (Definition 3.9) is nondeterministic in goal selection and abstracts away suspension and failure, dGLP provides:
  1. 1.
    Deterministic FIFO scheduling of active goals
  2. 2.
    Explicit tracking of suspended goals with their suspension sets
  3. 3.
    Explicit tracking of failed goals
  4. 4.
    Automatic reactivation of suspended goals when blocking readers are instantiated
The key simplification compared to multiagent GLP is that all variables are local, so reader assignments can be applied immediately rather than through asynchronous communication.
Definition 3.14 (Reduction Suspension and Failure).
Given unit goal $A$ and clause $C$, with $H$ :- $B$ being the result of the GLP renaming of $C$ apart from $A$:
  • The reduction succeeds if $A$ and $H$ have a writer mgu $\sigma$ (Definition 3.8)
  • The reduction suspends with suspension set $W$ if no writer mgu exists, but there exists a set $W\subseteq\mathcal{V}?$ of readers occurring in $A$ such that for some ground substitution $\theta$ on $W$, the terms $A\theta$ and $H$ have a writer mgu. $W$ is the minimal such set.
  • The reduction fails if it neither succeeds nor suspends
Definition 3.15 (dGLP Configuration).
A dGLP configuration over program $P$ is a triple $(Q,S,F)$ where:
  • $Q$ is a sequence (queue) of active unit goals
  • $S$ is a set of suspended unit goals, each paired with its suspension set of blocking readers
  • $F$ is a set of failed unit goals
Definition 3.16 (Reactivation Set).
Given a suspended goal set $S$ and a readers substitution $\sigma?$, the reactivation set is:
$\textrm{reactivate}(S,\sigma?)=\{G:(G,W)\in S\wedge\exists X?\in W.\;X?\sigma?\neq X?\}$
These are goals suspended on readers that have been instantiated by $\sigma?$.
Definition 3.17 (dGLP Transition System).
The transition system $\textrm{dGLP}(P)=(\mathcal{C},c_{0},\mathcal{T})$ over GLP program $P$ and initial goal $G_{0}$ satisfying SO is defined by:
  1. 1.
    $\mathcal{C}$ is the set of all dGLP configurations over $P$
  2. 2.
    $c_{0}=(G_{0},\emptyset,\emptyset)$
  3. 3.
    $\mathcal{T}$ is the set of all transitions $(Q,S,F)\rightarrow(Q^{\prime},S^{\prime},F^{\prime})$ where $Q=A\cdot Q_{r}$ (head goal $A$, remainder $Q_{r}$) and exactly one of the following holds:
    1. (a)
[…]
      $Q^{\prime}=(Q_{r}\cdot B\cdot R)\hat{\sigma}\hat{\sigma}?,\quad S^{\prime}=S\setminus\{(G,W):G\in R\},\quad F^{\prime}=F$
    2. (b)
      Suspend: No clause succeeds, but the GLP reduction of $A$ with at least one clause $C\in P$ suspends. Let $W=\bigcup_{C\in P}W_{C}$ where $W_{C}$ is the suspension set for clause $C$. Then:
      $Q^{\prime}=Q_{r},\quad S^{\prime}=S\cup\{(A,W)\},\quad F^{\prime}=F$
    3. (c)
      Fail: No clause succeeds and no clause suspends (all clauses fail outright). Then:
      $Q^{\prime}=Q_{r},\quad S^{\prime}=S,\quad F^{\prime}=F\cup\{A\}$
dGLP Correctly implements GLP.
Definition 3.18 (Implementation Mapping $\sigma$).
Given a dGLP computation $r^{\prime}=c^{\prime}_{0}\rightarrow c^{\prime}_{1}\rightarrow\ldots\rightarrow c^{\prime}_{n}$ where $c^{\prime}_{n}=(Q,S,F)$, the implementation mapping $\sigma$ is defined by:
$\sigma(c^{\prime}_{n})=(G,\sigma_{r})$
where $G=Q\cup\{A:(A,W)\in S\}\cup F$ is the multiset union of active, suspended, and failed goals, and $\sigma_{r}=\hat{\sigma}_{1}?\circ\hat{\sigma}_{2}?\circ\ldots\circ\hat{\sigma}_{k}?$ is the composition of all readers substitutions produced by Reduce transitions along $r^{\prime}$.
Lemma 3.19 ($\sigma$ is Live).
The implementation mapping $\sigma$ is live.
While dGLP cannot match GLP’s arbitrary reduction order step-by-step (GLP is nondeterministic in goal selection; dGLP uses FIFO), it achieves the same outcomes. This follows from the commutativity of disjoint substitution compositions.
Lemma 3.20 (Disjoint Substitution Commutativity).
Let $\hat{\sigma}_{1}$ and $\hat{\sigma}_{2}$ be writers substitutions produced by two GLP Reduce transitions on distinct goals $A_{1}$ and $A_{2}$ in the same run. Then $\hat{\sigma}_{1}$ and $\hat{\sigma}_{2}$ assign disjoint sets of variables, and $\hat{\sigma}_{1}\circ\hat{\sigma}_{2}=\hat{\sigma}_{2}\circ\hat{\sigma}_{1}$.
Lemma 3.21 (Reduction Set Determines Outcome).
Let $\rho_{1}$ and $\rho_{2}$ be two proper GLP runs from $(G_{0},\emptyset)$ that perform the same set of Reduce transitions $\{R_{1},\ldots,R_{k}\}$ (each $R_{i}$ being a reduction of some goal with some clause). Then $\rho_{1}$ and $\rho_{2}$ have the same outcome.
Lemma 3.22 (dGLP Performs Same Reductions).
For every proper GLP run $\rho$ that performs Reduce transitions $\{R_{1},\ldots,R_{k}\}$, there exists a dGLP run $\rho^{\prime}$ that performs the same set of Reduce transitions.
Lemma 3.23 (dGLP Completeness).
For every proper GLP run $\rho$ with outcome $O$, there exists a dGLP run $\rho^{\prime}$ with outcome $O$.
Theorem 3.24 (dGLP Correctly Implements GLP).
The implementation $(dGLP,\sigma)$ of GLP is correct.
Proof 3.25.
By Definition 2.3, correctness requires liveness and completeness.
Live: By Lemma 3.19.
Complete: By Lemma 3.23.
All lemma proofs and additional remarks are in Appendix A.
4 Multiagent Transition Systems
This section extends transition systems to the multiagent setting, using atomic transactions to specify multiagent behaviors.
4.1 Agents, Local States, and Configurations
We assume a potentially infinite set of agents $\Pi$, but consider only finite subsets of it, so when we refer to a particular set of agents $P\subset\Pi$ we assume $P$ to be nonempty and finite. We use $\subset$ to denote the strict subset relation and $\subseteq$ when equality is also possible.
We use $S^{P}$ to denote the set $S$ indexed by the set $P$, and if $c\in S^{P}$ we use $c_{p}$ to denote the member of $c$ indexed by $p\in P$.
Definition 4.1 (Local States, Configuration, Transaction, Participants).
Given agents $Q\subset\Pi$ and an arbitrary set $S$ of local states, a configuration over $Q$ and $S$ is a member of $C:=S^{Q}$. An atomic transaction, or just transaction, over $Q$ and $S$ is any pair of configurations $t=c\rightarrow c^{\prime}\in C^{2}$ such that $c\neq c^{\prime}$, with $t_{p}:=c_{p}\rightarrow c^{\prime}_{p}$ for any $p\in Q$, and with $p$ being an active participant in $t$ if $c_{p}\neq c^{\prime}_{p}$, stationary participant otherwise.
Definition 4.2 (Degree).
The degree of a transaction $t$ (unary, binary, …, $k$-ary) is the number of active participants in $t$, and the degree of a set of transactions $T$ is the maximal degree of any $t\in T$.
4.2 Multiagent Transition Systems
Definition 4.3 (Multiagent Transition System).
Given agents $P\subset\Pi$ and an arbitrary set $S$ of local states with a designated initial local state $s0\in S$, a multiagent transition system over $P$ and $S$ is a transition system $TS=(C,c0,T)$ with configurations $C:=S^{P}$, initial configuration $c0:=\{s0\}^{P}$, and transitions $T\subseteq C^{2}$ being a set of transactions over $P$ and $S$, with the degree of $TS$ being the degree of $T$.
Definition 4.4 (Transaction Closure).
Let $P\subset\Pi$, $S$ a set of local states, and $C:=S^{P}$. For a transaction $t=(c\rightarrow c^{\prime})$ over local states $S$ with participants $Q\subseteq P$, the $P$-closure of $t$, $t{\uparrow}P$, is the set of transitions over $P$ and $S$ defined by:
$t{\uparrow}P:=\{t^{\prime}\in C^{2}:\forall q\in Q.(t_{q}=t^{\prime}_{q})\wedge\forall p\in P\setminus Q.(p\text{ is stationary in }t^{\prime})\}$
If $R$ is a set of transactions, each $t\in R$ over some $Q\subseteq P$ and $S$, then the $P$-closure of $R$, $R{\uparrow}P$, is the set of transitions over $P$ and $S$ defined by:
$R{\uparrow}P:=\bigcup_{t\in R}t{\uparrow}P$
Definition 4.5 (Transactions-Based Multiagent Transition System).
Given agents $P\subset\Pi$, local states $S$ with initial local state $s0\in S$, and a set of transactions $R$, each $t\in R$ over some $Q\subseteq P$ and $S$, the transactions-based multiagent transition system over $P$, $S$, and $R$ is the multiagent transition system $TS=(S^{P},\{s0\}^{P},R{\uparrow}P)$.
5 Multiagent GLP and Its Implementation
5.1 Multiagent GLP
This section extends GLP to multiple agents, defining multiagent GLP (maGLP) as a transactions-based multiagent transition system using the framework developed in Section 4.
From GLP to Multiagent GLP. In extending GLP to multiple agents, each agent maintains its own asynchronous resolvent as its local state. The key insight is that GLP’s variable pairs provide natural binary communication channels: when agent $p$ assigns a writer $X$ for which the paired reader $X?$ is held by agent $q$, the assignment $X:=T$ must be communicated to $q$.
A key difference between single-agent GLP and multiagent GLP is in the initial state. In a multiagent transition system all agents must have the same initial local state $s0$ (Definition 4.3). This precludes setting up an initial configuration in which agents share logic variables, as this would imply different initial states for different agents.
We resolve this in two steps. First, we employ only anonymous logic variables “_” in the initial local states of agents: Anonymous variables are, on the one hand, syntactically identical, hence allow all initial states to be syntactically identical, and on the other hand represent unique variables, hence semantically all initial goals have unique, local, non-shared variables. The initial state of all agents is the atomic goal agent(ch(_?,_),ch(_?,_)), with the first channel serving communication with the user and the second with the network.
Second, the Cold-call transaction enables agents to bootstrap communication by establishing shared variables through the network infrastructure, realizing the cold-call protocol for connecting previously-disconnected agents.
Definition 5.1 (Multiagent GLP).
[…]
  1. 1.
    Reduce $p$: A unary transaction with participant $p$ where $c_{p}\rightarrow c^{\prime}_{p}$ is a GLP Reduce transition (Definition 3.9).
  2. 2.
    Communicate $p$ to $q$: A transaction with participants $p,q\in P$ where $c_{p}=(G_{p},\sigma_{p})$, $c_{q}=(G_{q},\sigma_{q})$, $\{X?:=T\}\in\sigma_{p}$, $X?$ occurs in $G_{q}$, $c^{\prime}_{p}=(G_{p},\sigma_{p}\setminus\{X?:=T\})$, and $c^{\prime}_{q}=(G_{q}\{X?:=T\},\sigma_{q})$.
  3. 3.
    Cold-call $p$ to $q$: A binary transaction with participants $p\neq q\in P$ where the network output stream in $c_{p}$ has a new message msg(q,X)(q,X), $c^{\prime}_{p}$ is the result of advancing the network output stream in $c_{p}$, and $c^{\prime}_{q}$ is the result of adding $X?$ to the network input stream in $c_{q}$.
Note that Communicate may be unary or binary, depending on whether $p=q$. Both Communicate and Cold-call transfer assignments from writers to readers: Communicate operates between agents sharing a paired reader and writer, while Cold-call operates through the network streams established in each agent’s initial configuration, enabling the creation of paired variables among previously-disconnected agents.
Lemma 5.2 (maGLP Persistence).
maGLP is persistent.
See Appendix B for proofs and further remarks.
Definition 5.3 (maGLP Fair Run).
A complete maGLP run is fair if every transaction that becomes enabled is eventually taken.
Definition 5.4 (maGLP Proper Run, Outcome).
An maGLP run is proper if the run at each agent is proper (Definition 3.10). The outcome of a proper maGLP run is the tuple of per-agent outcomes.
5.2 Implementing Multiagent GLP with Deterministic Agents
Motivation. madGLP serves as the formal specification for AI to implement multiagent GLP on smartphones communicating peer-to-peer, without reliance on central servers. The discipline is: from the mathematical definition, AI derives an English+code specification, and from the spec derives Dart code. The complete specification is in Appendix C.
This section specifies madGLP, a multiagent deterministic transition system that implements maGLP using only local variable pairs connected by global links. While maGLP defines shared variable pairs that span agent boundaries, madGLP replaces each such pair with two fully local pairs connected through a global writers table and message passing, with forwarding handled by spawned GLP goals.
Implementing Shared Variables via Local Pairs. The key insight of madGLP is that a maGLP shared variable pair $(X,X?)$ with writer $X$ at agent $p$ and reader $X?$ at agent $q$ can be implemented by two local pairs connected by a global link:
  • At agent $p$: a local pair $(X_{p},X_{p}?)$ where both variables remain in $p$’s resolvent
  • At agent $q$: a local pair $(X_{q},X_{q}?)$ where both variables remain in $q$’s resolvent
  • A global link connecting $X_{p}$ to $X_{q}$, realized as a global_send goal at $p$ that monitors $X_{p}?$ and an entry in $q$’s global writers table that maps the global name to $X_{q}$
When $X_{p}$ is assigned a term $T$ at agent $p$, the value becomes available via $X_{p}?$. A spawned global_send goal detects this and sends the assignment message to $q$, where $T$ is localized. Upon receipt, $q$ looks up the target writer in its global writers table, assigns $X_{q}:=T_{q}^{\downarrow}$, and removes the entry. This makes the value available through $X_{q}?$ in $q$’s resolvent.
This design keeps all variables in the resolvent and handles forwarding uniformly through GLP goals, enabling correct behavior even when both ends of a variable pair are exported to the same or different agents.
Global Names, Tables, and Communication. Global variable names _w(p,i)\_w(p,i) and _r(p,i)\_r(p,i) identify the endpoints of global links in messages (Definition C.1 in Appendix C). The global writers table at each agent tracks local writers awaiting incoming assignments (Definition C.3 in Appendix C). Outgoing communication is handled uniformly by spawned global_send goals (Definition C.5 in Appendix C).
Cold-calls—messages between agents with no prior shared variables—use a reserved index-0 serializer that allows multiple senders to append to an agent’s network input stream (Definition C.7 in Appendix C).
Globalizing and Localizing Terms. Globalization ($T^{\uparrow}$) converts local variables to global names for message passing; localization ($T^{\downarrow}$) converts global names back to fresh local pairs (Definitions C.12 and C.13 in Appendix C). The pairing ensures correct dataflow for both writers and readers, including the case where both ends of a variable pair are exported.
madGLP Local States and Transitions.
Definition 5.5 (madGLP Local State).
The local state of agent $p\in\Pi$ is a tuple $s_{p}=(R_{p},W_{p},M_{p})$ where:
  1. 1.
    $R_{p}=(A_{p},S_{p},F_{p})$ is a deterministic resolvent: $A_{p}$ is a queue of active goals, $S_{p}$ contains suspended goals paired with blocking readers, and $F_{p}$ contains failed goals
  2. 2.
    $W_{p}$ is a global writers table
  3. 3.
    $M_{p}$ is a set of pending outgoing messages
Definition 5.6 (madGLP Transition System).
The madGLP transition system over agents $P\subset\Pi$ and GLP program $M$ is the multiagent transition system $\textrm{madGLP}=(C,c0,T)$ where $C$ is the set of configurations, $c0$ is the initial configuration with empty resolvents except for the initial agent goal and index-0 serializer entry, and $T$ consists of:
  • Reduce: unary transaction processing goals via FIFO scheduling
  • Send: unary transaction moving messages to communication channels
  • Receive: unary transaction processing incoming messages
Full transaction definitions are in Appendix C.
All madGLP transactions are unary. Cold-calls, which in maGLP require binary transactions, are implemented using the index-0 serializer, decomposing into unary Send and Receive transactions. We assume fair message delivery, a standard assumption for distributed systems with reliable channels: every message placed in a communication channel is eventually delivered to its recipient.
Example: Client-Monitor Communication. We illustrate madGLP’s variable linking with a client-monitor scenario.
Program 1: Concurrent Monitor
monitor(Reqs) :- monitor(Reqs?,0).
monitor([add|Reqs],Sum) :-
    Sum1 := Sum? + 1, monitor(Reqs?,Sum1?).
monitor([subtract|Reqs],Sum) :-
    Sum1 := Sum? - 1, monitor(Reqs?,Sum1?).
monitor([value(Sum?)|Reqs],Sum) :-
    integer(Sum?) |  monitor(Reqs?,Sum?).
monitor([],_).
Consider the initial goal client1(Xs)@p, monitor(Xs?)@q, which establishes a shared pair with writer $\mathsf{Xs}$ at $p$ and reader $\mathsf{Xs}?$ at $q$. Figure 1 shows how maGLP shared variables correspond to madGLP local pairs connected by global links, through three stages:
  • Stage 0: At boot time, the coordinator globalizes the initial goal, creating local pairs at each agent connected by global links. Agent $p$ gets $\texttt{client1}(\mathsf{Xs_{p}})$ and spawns $\texttt{global\_send}(\mathsf{Xs_{p}}?,\_w(p,1),q)$; agent $q$ gets $\texttt{monitor}(\mathsf{Xs_{q}}?)$ and creates entry $(\mathsf{Xs_{q}},p,1)$.
  • Stage 1: After $p$ assigns $\mathsf{Xs}:=[\mathtt{add}|\mathsf{Xs1}]$, the global_send goal fires, sending the value to $q$. A new link is created for $\mathsf{Xs1}$ with its own global_send goal at $p$ and entry at $q$.
  • Stage 2: After $p$ assigns $\mathsf{Xs1}:=[\mathtt{value}(\mathsf{V?})|\mathsf{Xs2}]$. Note that $p$ sends reader $\mathsf{V?}$, so $p$ creates entry $(\mathsf{V},q)$ and $q$ spawns $\texttt{global\_send}(\mathsf{V_{q}}?,\_r(p,k),p)$—the dataflow points from $q$ to $p$.
[…]
Figure 1: maGLP vs madGLP variable linking through three stages. Left: maGLP shared variable pairs. Right: madGLP local pairs with global links (arrows from reader to writer, labeled with global names). Dashed lines separate agents $p$ and $q$. “gs” denotes global_send goals watching readers. In Stage 2, the V link arrow points from $q$ to $p$ (callback): when $q$ assigns $\mathsf{V_{q}}$, the goal gs($\mathsf{V_{q}?}$) sends the value back to $p$.
Correctness. Global links implement maGLP shared variable pairs via message passing. The implementation mapping $\pi$ from madGLP configurations to maGLP configurations reconstructs the resolvent and readers substitution while forgetting implementation details (goal classification, global writers tables, message queues, global_send goals). Full definitions are in Appendix C.
Theorem 5.7 (madGLP Correctly Implements maGLP).
The implementation $(\textrm{madGLP},\pi)$ of maGLP is correct.
Proof 5.8.
Live: By Lemma C.41.
Complete: By Lemma C.45.
Corollary 5.9 (Global Links Correctly Implement Shared Variables).
Local variable pairs connected by global links and global_send goals in madGLP correctly implement shared variable pairs in maGLP, including the case where both ends of a pair are exported.
Proof 5.10.
Follows from Theorem 5.7 together with Lemmas C.28 and C.30.
Full definitions and lemma proofs are in Appendix C.
6 Related Work
Concurrent Constraint Programming. GLP’s monotonic semantics builds directly on the concurrent constraint programming (CCP) paradigm established by Saraswat, Rinard, and Panangaden [saraswat1991semantic]. Their foundational insight—that concurrent agents communicate through a shared constraint store that grows monotonically—provides the semantic template for GLP’s single-assignment variables. In CCP, computation proceeds by adding constraints; constraints are never retracted. GLP inherits this monotonicity: variable assignments accumulate but are never undone, and agent states grow through message reception. The lattice-based constraint stores with monotonic information accumulation in CCP are the direct theoretical ancestor to GLP’s monotonic multiagent framework.
GLP belongs to the family of concurrent logic programming languages that emerged in the 1980s: Concurrent Prolog [shapiro1983subset], GHC [ueda1986guarded], and PARLOG [clark1986parlog]. These languages interpret goals as concurrent processes communicating through shared logical variables, using committed-choice execution with guarded clauses. A comprehensive survey [shapiro1989family] documents this family and its design space, detailing operational semantics of guarded Horn clause languages, single-assignment variables, and writer/reader communication mechanisms that GLP refines and extends.
A key evolution was flattening: restricting guards to primitive tests only. Flat Concurrent Prolog [mierowsky1985fcp] and Flat GHC demonstrated that flat guards suffice for practical programming while dramatically simplifying semantics and implementation. GLP can be understood as Flat Concurrent Prolog with the Single-Reader Single-Writer (SRSW) constraint added. This resolves the semantic difficulties of read-only unification identified by Levi and Palamidessi [levi1985readonly], ensuring that term matching suffices without full unification.
Modes and Linearity. Mode systems for concurrent logic programming assign polarity to variable occurrences. Ueda’s moded Flat GHC [ueda1994moded, ueda1995io] is most directly relevant: his mode system assigns polarity to every variable occurrence (positive for input/read, negative for output/write), with the well-modedness property guaranteeing each variable is written exactly once. This establishes the “programming as wiring” paradigm where programs describe dataflow networks. Ueda’s subsequent linearity analysis [ueda2001resource] identifies variables read exactly once, enabling compile-time garbage collection for single-reader structures.
GLP enforces both single-reader and single-writer universally as a syntactic restriction, whereas Ueda’s system guarantees single-writer with single-reader as an optional refinement. This stronger restriction enables GLP’s simpler distributed execution model.
The SRSW discipline connects to linear logic’s [girard1987linear, wadler1993taste] resource sensitivity: each variable assignment is produced at most once and consumed at most once. Hodas and Miller’s Lolli [hodas1994lolli] presents logic programming in a fragment of intuitionistic linear logic where resources must be used exactly once unless marked with the exponential modality—conceptually related to GLP’s variable restrictions.
Session Types and Linear Logic. GLP’s writer/reader pairs have deep connections to session types via the Curry-Howard correspondence for concurrency. Caires and Pfenning [caires2010session] established this correspondence, interpreting linear logic propositions as session types where channels are used exactly once—directly paralleling GLP’s single-use semantics. Their treatment of dual session endpoints mirrors GLP’s writer/reader pair structure.
[…]
Honda, Yoshida, and Carbone’s multiparty asynchronous session types [honda2016multiparty] introduce global types and projection mechanisms for coordinating multiple agents—relevant to GLP’s multiagent setting where agents communicate through shared variables. A companion paper [shapiro2026types] develops a full type system for GLP based on these connections.
Futures and Promises. A GLP reader/writer pair is operationally a future or promise [baker1977future, friedman1976impact]: the writer creates a promise for a value, the reader awaits fulfillment. Flanagan and Felleisen [flanagan1999semantics] provide increasingly intensional operational semantics for futures, with their third semantics explicitly modeling placeholder objects and touch operations—the synchronization between value producers and consumers directly analogous to GLP’s single-assignment variables that block readers until writers provide values.
Most relevant is Pruiksma and Pfenning’s work on linear futures [pruiksma2022futures], developing concurrent languages based on adjoint logic that encode write-once shared-memory semantics analogous to single-assignment variables. They establish connections between futures and session types through Curry-Howard correspondence. Unlike futures in mainstream languages, GLP futures can carry logical terms containing further reader/writer pairs, enabling recursive communication structures—a capability shared with channel-passing calculi such as the $\pi$-calculus [milner1999communicating].
Transition Systems and Refinement. The correctness framework in this paper employs simulation relations between transition systems, following the I/O automata methodology of Lynch and Tuttle [lynch1987hierarchical, lynch1988introduction]. Their forward and backward simulation techniques [lynch1995forward, lynch1996forward] provide the template for proving that implementations correctly refine specifications: if every implementation trace is a valid specification trace, the implementation is correct.
Abadi and Lamport’s theory of refinement mappings [abadi1991existence] establishes that such mappings can always be found (under suitable conditions) by adding history and prophecy variables—a result awarded the LICS Test of Time Award for its foundational importance. Lamport’s TLA+ [lamport1994temporal] provides a related framework combining temporal logic with refinement. The refinement calculus [back1998refinement] and data refinement theory [deroever1998data] provide systematic frameworks for proving implementation correctness through stepwise refinement. Our proofs use forward simulation directly without requiring auxiliary variables.
Sangiorgi’s work on bisimulation proof methods [sangiorgi1998bisimulation] develops “up-to” techniques that reduce the size of relations needed to establish bisimilarity. While we use simulation (trace inclusion) rather than bisimulation (trace equivalence), similar proof techniques apply. The extension by Madiot, Pous, and Sangiorgi [madiot2014bisimulations] transports these techniques to sophisticated labeled transition systems including the $\pi$-calculus—applicable to GLP’s multiagent transition systems with term-passing communication.
Atomic Transactions. Our key technical contribution—that binary atomic transactions can be correctly implemented by sequences of unary transactions—builds on the theory of atomic transactions in distributed systems. Lynch et al.’s foundational treatment [lynch1988atomic, lynch1994atomic] establishes the formal framework for transaction serializability and correctness. Bernstein, Hadzilacos, and Goodman [bernstein1987concurrency] provide comprehensive coverage of concurrency control.
Herlihy and Wing’s linearizability [herlihy1990linearizability] defines the gold standard for concurrent object correctness: operations appear to take effect atomically at some point between invocation and response. Our transaction decomposition achieves a related goal—showing that multi-step implementations are equivalent to atomic specifications—using disjoint substitution commutativity (from the SO invariant) and Persistence (Lemma 3.12) rather than linearization points.
Grassroots and Monotonic Distributed Systems. The grassroots property—that any subset of agents forms a functioning subsystem without requiring permission or coordination—was introduced for distributed platforms [shapiro2023grassrootsBA] using multiagent transition systems with monotonicity as a key property. Applications include grassroots social networks [shapiro2023gsn], cryptocurrencies [shapiro2024gc, lewis2023grassroots], and federated governance [halpern2024federated, shapiro2025GF]. This paper provides general theorems (Appendix D) showing that correct and complete implementations preserve the grassroots property, enabling modular reasoning about grassroots systems at multiple levels of abstraction.
Conflict-free Replicated Data Types (CRDTs) [shapiro2011conflict] achieve eventual consistency without coordination by requiring that concurrent updates commute. The mathematical structure is similar: CRDTs use join-semilattices where states merge monotonically, just as GLP agents’ states grow monotonically through message reception. Kleppmann’s work on Byzantine-tolerant CRDTs [kleppmann2022making] extends this to adversarial settings. The Blocklace [almeida2024blocklace] demonstrates how independent subsystems can operate and later merge state—precisely the compositionality that grassroots systems require.
Related Language Implementations. Several concurrent languages share design goals with GLP. Erlang [armstrong2010erlang] provides lightweight processes with message passing; Dialyzer [lindahl2006practical] adds success typing. Go [pike2012go] offers goroutines and channels. Rust [matsakis2014rust] enforces ownership for memory safety. Session-typed languages like Links [cooper2006links, lindley2017lightweight] and SILL [toninho2013higher] integrate linear types with communication. Actor languages like Pony [clebsch2015deny] use capability systems for data-race freedom. The ABS language [johnsen2011abs] combines active objects with cooperative scheduling and futures.
GLP differs in its logic programming foundation: programs are declarative specifications that also serve as executable distributed code, and the SRSW discipline emerges from logical structure rather than a separate type system.
7 Conclusion
Development History. The madGLP design underwent two major revisions driven by implementation difficulties discovered during Dart programming. The initial design used routing tables to track variable locations; when variables migrated between agents, these tables required complex updates that proved error-prone to implement. The current design eliminates variable migration entirely: globally-shared variable pairs are realized by two local pairs connected by a global link from the writer of one to the reader of the other, with all variables remaining in their originating agent’s resolvent. Similarly, cold calls—messages between agents with no prior shared variables—were initially a separate mechanism; they are now realized uniformly as regular terms sent to a reserved index-0 shared variable that serializes incoming network messages.
Summary. We have presented dGLP and madGLP—implementation-ready operational semantics for single-agent and multiagent GLP respectively—and proved them correct with respect to their abstract counterparts. The key insight is that disjoint substitution commutativity (from the SO invariant) and Persistence (Lemma 3.12) enable correctness proofs that carry over directly from the single-agent to the multiagent case. The madGLP design—local variable pairs connected by global links—provides a clean separation between local computation and inter-agent communication, supporting both the formal correctness proofs and practical implementation.
Appendix A Single-Agent GLP Details
This appendix contains proofs and remarks for single-agent GLP (Section 3).
A.1 GLP Remarks
Remark A.1 (GLP is Persistent).
GLP is persistent: all transitions, once enabled, remain enabled until taken.
  • Reduce: By Persistence (Lemma 3.12) and the SO invariant, an enabled Reduce for goal $A$ (tracking $A$’s identity across configurations) remains enabled until taken, since Reduce transitions for other goals $B\neq A$ assign disjoint writers (by SO), and Communicate transitions apply only reader substitutions which do not affect writer mgu existence.
  • Communicate: An enabled Communicate applying $\{X?:=T\}$ to $X?\in G$ remains enabled, since Reduce transitions do not remove assignments from $\sigma$, and other Communicate transitions apply different assignments.
Remark A.2 (Properness).
The properness condition ensures fresh variables introduced by clause renaming are indeed fresh with respect to the entire computation history. This is guaranteed by the GLP renaming-apart requirement in Definition 3.6.
A.2 dGLP Proofs
See 3.12
Proof A.3.
For Reduce: GLP Reduce is enabled for $A$ iff there exists a clause $C$ with head $H$ such that $A$ and $H$ have a writer mgu $\hat{\sigma}$. By Definition 3.7, a writer mgu assigns only writers, not readers. By the SO restriction, each writer occurs at most once in $A$; by SRSW, each writer in $A$ is paired with a reader in $A$. Since $\sigma^{\prime}_{r}$ is a readers substitution (assigning only readers), $A\sigma^{\prime}_{r}$ and $H$ still have the same writer mgu $\hat{\sigma}$. Other Reduce transitions assign disjoint writers (by SO). Thus the Reduce remains enabled.
For Communicate: applying $\{X?:=T\}\in\sigma$ to $X?\in G$ remains enabled since Reduce transitions do not remove assignments from $\sigma$ and other Communicate transitions apply different assignments.
See 3.19
Proof A.4.
Since dGLP is deterministic (Remark A.9), every complete dGLP run is trivially fair—there is at most one enabled transition at any configuration, and it is taken. By Remark A.1, GLP is persistent.
Reduce: Since dGLP’s FIFO scheduling eventually processes every goal that enters the queue, every enabled GLP Reduce is eventually taken in $\sigma(r^{\prime})$.
Communicate: In dGLP, reader substitutions $\hat{\sigma}$? are applied immediately during Reduce (Remark A.10). Each such application corresponds to the GLP Communicate transitions that would apply those reader assignments. Hence every enabled GLP Communicate is implicitly taken when the corresponding dGLP Reduce fires.
Therefore $\sigma$ maps fair dGLP runs to fair GLP runs.
See 3.20
Proof A.5.
By the SO invariant, each writer occurs at most once in the goal multiset $G$. Since $A_{1}\neq A_{2}$ are distinct goals, they have disjoint writers. Each Reduce assigns only writers from its reduced goal (by Definition 3.7). Therefore $\hat{\sigma}_{1}$ and $\hat{\sigma}_{2}$ have disjoint domains. Substitutions with disjoint domains commute.
See 3.21
Proof A.6.
Let $\rho_{1}$ perform reductions in order $R_{\pi(1)},\ldots,R_{\pi(k)}$ and $\rho_{2}$ in order $R_{\tau(1)},\ldots,R_{\tau(k)}$ for permutations $\pi,\tau$.
Each $R_{i}$ produces a writers substitution $\hat{\sigma}_{i}$. By Lemma 3.20, these substitutions have disjoint domains and commute. Therefore:
$\hat{\sigma}_{\pi(1)}\circ\cdots\circ\hat{\sigma}_{\pi(k)}=\hat{\sigma}_{\tau(1)}\circ\cdots\circ\hat{\sigma}_{\tau(k)}$
The final readers substitution $\sigma_{n}$ is the composition of the readers counterparts, which likewise commute.
The final goal $G_{n}$ consists of goals added by clause bodies (which depends only on which reductions occurred, not their order) with the accumulated substitution applied (which is the same by the above).
Communicate transitions apply reader assignments but do not affect the outcome: they do not change $\sigma_{n}$ and merely propagate existing assignments within goals. Therefore both runs have the same outcome $(G_{0}$ :- $G_{n})\sigma_{n}$.
See 3.22
Proof A.7.
By induction on spawn order. Goals in the initial goal $G_{0}$ are in the initial dGLP queue. For any $R_{i}$ reducing $A_{i}$ with $C_{i}$: if $A_{i}$ was spawned by an earlier reduction $R_{j}$, by induction dGLP performs $R_{j}$, so $A_{i}$ enters the queue. Since $R_{i}$ succeeds in $\rho$, $A_{i}$ and the head of $C_{i}$ have a writer mgu. By Persistence (Lemma 3.12), this mgu exists regardless of readers substitution. When $A_{i}$ reaches the queue head, the reduction succeeds.
See 3.23
Proof A.8.
Let $\rho$ be a proper GLP run with outcome $O=(G_{0}$ :- $G_{n})\sigma_{n}$.
Let $\rho$ perform Reduce transitions $\{R_{1},\ldots,R_{k}\}$.
By Lemma 3.22, there exists a dGLP run $\rho^{\prime}$ that performs the same set of reductions $\{R_{1},\ldots,R_{k}\}$, possibly in a different order and with interleaved Suspend/reactivation steps.
By Lemma 3.21, since $\rho$ and $\rho^{\prime}$ perform the same reductions, they have the same outcome $O$.
A.3 dGLP Remarks
Remark A.9 (Deterministic Scheduling).
The dGLP transition system is deterministic: given any non-terminal configuration, exactly one transition is enabled. Goal selection follows FIFO order from the active queue $Q$, and clause selection uses the first applicable clause as in GLP.
Remark A.10 (Immediate Reader Application).
In dGLP, reader substitutions $\hat{\sigma}$? are applied immediately during reduction rather than through separate Communicate transitions. This is sound for single-agent execution where all variables are local.
Remark A.11 (Step-by-Step vs. Outcome Equivalence).
Step-by-step completeness would require dGLP to match GLP’s transitions in order, which is impossible due to GLP’s nondeterministic goal selection vs. dGLP’s FIFO scheduling. Outcome-based completeness is the appropriate notion: what matters for program correctness is the final result (the clause derived and final substitution), not the intermediate reduction order.
Appendix B Multiagent GLP Details
This appendix contains proofs and remarks for multiagent GLP (Section 5).
B.1 maGLP Proofs
See 5.2
Proof B.1.
maGLP Reduce is enabled for $A\in G_{p}$ iff there exists a clause $C$ with head $H$ such that $A$ and $H$ have a writer mgu $\hat{\sigma}$. Communicate applies a readers substitution $\{X?:=T\}$ to $G_{p}$; by Definition 3.7, writer mgu assigns only writers, so $A\{X?:=T\}$ and $H$ still have the same writer mgu $\hat{\sigma}$. Cold-call adds readers to $G_{q}$ but does not modify $G_{p}$’s existing goals. In both cases, the Reduce remains enabled.
B.2 maGLP Remarks
Remark B.2 (Transaction Degrees).
Reduce is unary; Communicate and Cold-call are binary (unary when $p=q$).
Remark B.3 (maGLP is Persistent).
maGLP is persistent: all transactions, once enabled, remain enabled until taken.
  • Reduce: By Persistence (Lemma 3.12) and the SO invariant, an enabled Reduce for goal $A$ at agent $p$ (tracking $A$’s identity across configurations) remains enabled until taken, since Reduce transactions for other goals assign disjoint writers (by SO), Communicate transactions apply only reader substitutions, and Cold-call transactions only add new goals.
  • Communicate: An enabled Communicate from $p$ to $q$ remains enabled, since other transactions do not remove reader assignments from the readers substitution.
  • Cold-call: An enabled Cold-call from $p$ to $q$ remains enabled, since other transactions do not consume the network output stream message.
Appendix C madGLP Specification
This appendix contains the complete madGLP specification: full definitions, remarks, and correctness proofs. It serves as the formal specification from which AI derives the English+code spec and Dart implementation.
C.1 Global Variable Names
Definition C.1 (Global Variable Name).
A global variable name is a term of the form _w(p,i)\_w(p,i) or _r(p,i)\_r(p,i), where $p\in\Pi$ is an agent identifier and $i\in\mathbb{N}$ is an index allocated by $p$ during globalization. The name _w(p,i)\_w(p,i) denotes a writer globalized at $p$, and _r(p,i)\_r(p,i) denotes a reader globalized at $p$.
Global variable names appear only in messages between agents, never in resolvents. They identify the source of a global link and enable message routing.
C.2 Global Writers Table
The global writers table tracks local writers that await incoming assignments from remote agents. Each entry maps a global name to the local writer that will be assigned when a message arrives.
Definition C.2 (Global Writers Table Entry).
A global writers table entry at agent $p$ is either:
  • $(X,q)$ for entries created by Globalize: $X\in\mathcal{V}$ is a local writer that will be assigned when a message arrives from agent $q$
  • $(X,q,i)$ for entries created by Localize: $X\in\mathcal{V}$ is a local writer, $q\in\Pi$ is the remote agent, and $i\in\mathbb{N}$ is the index in $q$’s global name (needed to match incoming messages)
Definition C.3 (Global Writers Table).
The global writers table $W_{p}$ of agent $p$ is an array of global writers table entries. For entries created by Globalize at index $i$, the index $i$ is the index in the global name _r(p,i)\_r(p,i). For entries created by Localize, the entry stores the remote index explicitly.
Remark C.4 (What the Global Writers Table Stores).
The table contains only writers that await incoming assignments. When agent $p$ globalizes a reader $Y?$, it creates an entry for writer $Y$ because $p$ will receive the assignment. When agent $q$ localizes a writer global name _w(p,i)\_w(p,i), it creates an entry for its local writer because $q$ will receive the assignment. No entries are created for outgoing links—those are handled by global_send goals.
C.3 The global_send Predicate
Outgoing communication is handled by spawned goals rather than by table entries. This uniform approach correctly handles all cases, including forwarding when both ends of a variable pair are exported.
Definition C.5 (global_send Predicate).
The system predicate global_send/3 is defined as:
global_send(T, G, Q) :- known(T) | ’_send’(T, G, Q).
where:
  • $T$ is the reader whose value will be sent when known
  • $G$ is the global variable name (_w(p,i)\_w(p,i) or _r(p,i)\_r(p,i)) identifying the link
  • $Q$ is the destination agent
The guard known(T) succeeds when $T$ is assigned a non-variable term. The builtin ’_send’(T, G, Q) globalizes $T$ and adds message $(G:=T^{\uparrow},Q)$ to the agent’s outgoing message set.
Remark C.6 (Forwarding via global_send).
When an agent exports both ends of a variable pair (e.g., sending $[X,X?]$ to another agent, or $X$ to one agent and $X?$ to another), the Globalize operation spawns a global_send goal for the exported writer. If a value arrives on one global link and is assigned to a local writer $X$, then $X?$ becomes known, triggering any global_send goal watching $X?$. This automatically forwards the value without requiring special forwarding logic in the Receive transaction.
C.4 Index-0 Serializer for Cold-Calls
Cold-calls—messages between agents with no prior shared variables—are implemented using a reserved global name at index 0. Each agent $p$ has a well-known global reader _r(p,0)\_r(p,0) that serves as a serializer for incoming cold-call messages.
Definition C.7 (Index-0 Serializer).
Index 0 in each agent’s global writers table is reserved for the network input serializer:
  • At boot time, agent $p$ creates a permanent entry $(N_{p},*)$ at index 0, where $N_{p}$ is the local writer for $p$’s network input stream
  • The global name _r(p,0)\_r(p,0) is the serializer reader; any agent can send to _w(p,0)\_w(p,0)
  • Index allocation for regular global links starts at 1
Remark C.8 (Cold-Call via Serializer).
To send a cold-call message $T$ to agent $q$, agent $p$ spawns the goal $\texttt{global\_send}(T,\_w(q,0),q)$. When $T$ becomes known, the ’_send’ builtin detects that index is 0 and uses serializer semantics:
  1. 1.
    The message sent is $w(q,0):=[T^{\uparrow}\mid\_w(q,0)]$, wrapping $T$ in a list cell with the serializer writer as tail
  2. 2.
    When $q$ receives this message, it assigns $N_{q}:=[T_{q}^{\downarrow}\mid N^{\prime}_{q}]$ where $N^{\prime}_{q}$ is a fresh local writer
  3. 3.
    The entry at index 0 is updated to $(N^{\prime}_{q},*)$, not removed
This allows multiple agents to send cold-calls to $q$ concurrently; messages are serialized onto $q$’s network input stream in an arbitrary order determined by message arrival.
Remark C.9 (Network Output Processing).
Each agent spawns a send_to_net goal to process its network output stream:
send_to_net([msg(Q, T) | In]) :-
    global_send(T?, _w(Q,0), Q?), send_to_net(In?).
send_to_net([]).
When the agent writes msg(q, T) to its network output stream, the send_to_net goal observes this and spawns $\texttt{global\_send}(T?,\_w(q,0),q)$. This uses the standard global_send/3 mechanism with the well-known serializer address at index 0, requiring no special treatment beyond the index-0 semantics already defined.
Remark C.10 (Serializer vs Regular Global Links).
For index $i>0$, the ’_send’ builtin sends $G:=T^{\uparrow}$ directly and the receiver removes the entry after assignment. For index 0 (serializer), the builtin wraps the term in a list cell $T^{\uparrow}\mid G]$ reusing the serializer writer, and the entry is updated rather than removed. This distinction is the only special case for cold-calls; all other machinery (globalization, localization, message routing) is shared with regular global links.
Remark C.11 (Potential Generalization).
The many-to-one serializer pattern at index 0 implements an efficient multiway merge for network input. This pattern could be generalized as a language extension allowing programs to create additional serializers, but we leave such extensions to future work.
C.5 Globalizing and Localizing Terms
Definition C.12 (Globalize).
Given agent $p$, remote agent $q$, and term $T$, the globalization by $p$, written $T_{p}^{\uparrow}$, may update the global writers table $W^{\prime}_{p}$ and spawn goals into $p$’s resolvent as follows. For each variable $Y$ occurring in $T$:
  1. 1.
    If $Y$ is a writer: allocate the next index $i$, replace $Y$ in $T_{p}^{\uparrow}$ with _w(p,i)\_w(p,i), and spawn goal $\texttt{global\_send}(Y?,\_w(p,i),q)$ into $p$’s resolvent. No entry is created—the global_send goal handles outgoing communication.
  2. 2.
    If $Y?$ is a reader: allocate the next index $i$, create entry $(Y,q)$ at index $i$ in $W^{\prime}_{p}$, and replace $Y?$ in $T_{p}^{\uparrow}$ with _r(p,i)\_r(p,i). No goal is spawned—$p$ will receive the assignment on this link.
Definition C.13 (Localize).
Given agent $q$, remote agent $p$, and globalized term $T_{p}^{\uparrow}$, the localization by $q$, written $T_{q}^{\downarrow}$, may update the global writers table $W^{\prime}_{q}$ and spawn goals into $q$’s resolvent as follows. For each global name in $T_{p}^{\uparrow}$:
  1. 1.
    If _w(p,i)\_w(p,i): create fresh local pair $(Y_{q},Y_{q}?)$, allocate the next index $k$ in $W^{\prime}_{q}$, add entry $(Y_{q},p,i)$, and replace _w(p,i)\_w(p,i) with $Y_{q}?$ (the reader) in $T_{q}^{\downarrow}$. No goal is spawned—$q$ will receive the assignment on this link.
  2. 2.
    If _r(p,i)\_r(p,i): create fresh local pair $(Z_{q},Z_{q}?)$, replace _r(p,i)\_r(p,i) with $Z_{q}$ (the writer) in $T_{q}^{\downarrow}$, and spawn goal $\texttt{global\_send}(Z_{q}?,\_r(p,i),p)$ into $q$’s resolvent. No entry is created—the global_send goal handles outgoing communication.
Remark C.14 (Globalize-Localize Correspondence).
The pairing between Globalize and Localize ensures correct dataflow:
  • Writer globalized at $p$: Globalize spawns $\texttt{global\_send}(Y?,\_w(p,i),q)$ at $p$. Localize adds entry $(Y_{q},p,i)$ at $q$ and puts $Y_{q}?$ in $q$’s term. When $p$ assigns $Y$, the spawned goal fires and sends the value to $q$, where the entry routes it to $Y_{q}$, making it available via $Y_{q}?$.
  • Reader globalized at $p$: Globalize adds entry $(Y,q)$ at $p$. Localize puts $Z_{q}$ in $q$’s term and spawns $\texttt{global\_send}(Z_{q}?,\_r(p,i),p)$. When $q$ assigns $Z_{q}$, the spawned goal fires and sends the value to $p$, where the entry routes it to $Y$, making it available via $Y?$.
Remark C.15 (Exporting Both Ends of a Pair).
Consider agent $p$ exporting term $[X,X?]$ to agent $q$. Globalize processes both (indices start at 1 since 0 is reserved for the serializer):
  • Writer $X$: spawns $\texttt{global\_send}(X?,\_w(p,1),q)$, no entry
  • Reader $X?$: entry $(X,q)$ at index 2, no spawn
At $q$, Localize creates two independent pairs:
  • For _w(p,1)\_w(p,1): pair $(Y_{q},Y_{q}?)$, entry $(Y_{q},p,1)$, term gets $Y_{q}?$
  • For _r(p,2)\_r(p,2): pair $(Z_{q},Z_{q}?)$, term gets $Z_{q}$, spawns $\texttt{global\_send}(Z_{q}?,\_r(p,2),p)$
The term at $q$ is $Y_{q}?,Z_{q}]$. When $q$ assigns $Z_{q}:=T$:
  1. 1.
    $Z_{q}?$ becomes known (= $T$)
  2. 2.
    $\texttt{global\_send}(Z_{q}?,\_r(p,2),p)$ fires, sends $r(p,2):=T^{\uparrow}$ to $p$
  3. 3.
    $p$ receives, finds entry $(X,q)$ at index 2, assigns $X:=T^{\downarrow}$
  4. 4.
    $X?$ becomes known (= $T$)
  5. 5.
    $\texttt{global\_send}(X?,\_w(p,1),q)$ fires, sends $w(p,1):=T^{\uparrow}$ to $q$
  6. 6.
    $q$ receives, finds entry $(Y_{q},p,1)$, assigns $Y_{q}:=T^{\downarrow}$
  7. 7.
    $Y_{q}?$ becomes known (= $T$)
The value flows from $Z_{q}$ through $p$’s local pair to $Y_{q}?$, correctly implementing the semantics where both ends of the exported pair eventually share the same value.
C.6 madGLP Transactions
Definition C.16 (madGLP Reduce Transaction).
The unary Reduce transaction for agent $p$ transitions $s_{p}\rightarrow s^{\prime}_{p}$ where $A_{p}=A\cdot A_{r}$:
[…]
  1. 1.
    Apply $\hat{\sigma}$ to assign writers and $\hat{\sigma}$? to propagate to readers in the resolvent
  2. 2.
    Let $R$ be goals in $S_{p}$ suspended on readers that were instantiated; move them to active
  3. 3.
    $A^{\prime}_{p}:=(A_{r}\cdot B\cdot R)\hat{\sigma}\hat{\sigma}$?, update $S^{\prime}_{p}$ accordingly
Case Suspend: No clause succeeds but suspension set $W\neq\emptyset$: $A^{\prime}_{p}:=A_{r}$, $S^{\prime}_{p}:=S_{p}\cup\{(A,W)\}$
Case Fail: No clause succeeds and $W=\emptyset$: $A^{\prime}_{p}:=A_{r}$, $F^{\prime}_{p}:=F_{p}\cup\{A\}$
Definition C.17 (madGLP Send Transaction).
The unary Send transaction for agent $p$ is enabled when $(m,q)\in M_{p}$. It removes $(m,q)$ from $M_{p}$ and places $m$ in the communication channel to $q$.
Definition C.18 (madGLP Receive Transaction).
The unary Receive transaction processes a message $m$ from the communication channel:
Case $m=(\_w(q,0):=[T^{\uparrow}\mid\_w(q,0)])$ (serializer): Agent $q$ finds entry $(N_{q},*)$ at index 0. Localize $T^{\uparrow}$ by $q$ to get $T_{q}^{\downarrow}$, create fresh local writer $N^{\prime}_{q}$, assign $N_{q}:=[T_{q}^{\downarrow}\mid N^{\prime}_{q}?]$, apply $\{N_{q}?:=[T_{q}^{\downarrow}\mid N^{\prime}_{q}?]\}$ to goals, reactivate suspended goals, and update the entry to $(N^{\prime}_{q},*)$.
Case $m=(\_w(p,i):=T^{\uparrow})$ where $i>0$: The message is destined for the agent that localized _w(p,i)\_w(p,i). Agent $q$ searches its global writers table for an entry $(X_{q},p,i)$ matching the remote agent $p$ and remote index $i$. Localize $T^{\uparrow}$ by $q$ to get $T_{q}^{\downarrow}$, assign $X_{q}:=T_{q}^{\downarrow}$, apply $\{X_{q}?:=T_{q}^{\downarrow}\}$ to goals containing $X_{q}?$, reactivate suspended goals, and remove the entry from $W^{\prime}_{q}$.
Case $m=(\_r(p,i):=T^{\uparrow})$: The message is destined for agent $p$ who created this global name. Agent $p$ finds entry $(X,q)$ at index $i$ in $W_{p}$. Localize $T^{\uparrow}$ by $p$ to get $T_{p}^{\downarrow}$, assign $X:=T_{p}^{\downarrow}$, apply $\{X?:=T_{p}^{\downarrow}\}$ to goals containing $X?$, reactivate suspended goals, and remove the entry from $W^{\prime}_{p}$.
C.7 madGLP Remarks
Remark C.19 (Outgoing Messages via global_send).
The Reduce transaction does not directly generate outgoing messages. Instead, when a writer $X$ is assigned, $X?$ becomes known, which may trigger a global_send goal watching $X?$. That goal’s reduction (via the ’_send’ builtin) adds the message to $M_{p}$. This uniform approach handles both direct sends and forwarding.
Remark C.20 (Automatic Forwarding).
The Receive transaction simply assigns the local writer and removes the entry. If the assigned writer’s reader ($X?$) is being watched by a global_send goal (because it was also exported), that goal will fire on a subsequent Reduce, automatically forwarding the value. No special forwarding logic is needed in Receive.
Remark C.21 (Transaction Degrees).
All madGLP transactions are unary: Reduce operates on a single agent’s resolvent, Send removes a message from one agent’s outbox, and Receive processes an incoming message at one agent. Cold-calls, which in maGLP require binary transactions, are implemented in madGLP using the index-0 serializer mechanism, which decomposes into unary Send and Receive transactions.
Remark C.22 (Determinism).
madGLP is deterministic at the agent level: given a local state, exactly one Reduce transition is enabled (FIFO goal selection, first-clause matching). However, madGLP is nondeterministic at the system level due to communication asynchrony: the order in which messages arrive at an agent depends on network timing, and different arrival orders may lead to different computations.
Remark C.23 (Correspondence to maGLP).
The maGLP transactions are implemented in madGLP as follows:
  • Communicate: The maGLP binary Communicate transaction, which atomically transfers an assignment from one agent’s writer to another agent’s reader, is implemented by the sequence: Reduce (assigns writer, triggering global_send) $\rightarrow$ Send $\rightarrow$ Receive (applies assignment).
  • Cold-call: The maGLP binary Cold-call transaction is implemented using the index-0 serializer: the sender spawns $\texttt{global\_send}(T,\_w(q,0),q)$, which fires when $T$ is known, followed by Send $\rightarrow$ Receive (serializer appends to $q$’s network input stream).
The correctness of this implementation is proved below.
Remark C.24 (Global Writers Table Lifecycle).
An entry is added to the global writers table when a global link is established with the agent as the receiver: either when globalizing a reader (expecting an incoming assignment) or when localizing a writer global name (expecting the assignment). The entry is removed when the assignment arrives and the writer is assigned. The table tracks only those writers awaiting incoming assignments from remote agents.
Remark C.25 (Message Routing).
Messages use global names to enable routing:
  • Message _w(p,i):=T\_w(p,i):=T is sent from $p$ to whoever localized _w(p,i)\_w(p,i); that agent searches for an entry with remote agent $p$ and remote index $i$
  • Message _r(p,i):=T\_r(p,i):=T is sent to $p$ (the original globalizer); $p$ has an entry at index $i$ in its global writers table
C.8 Variable Correspondence
Definition C.26 (Global Link).
A global link from agent $p$ to agent $q$ for a writer consists of:
  • At agent $p$: a local variable pair $(X_{p},X_{p}?)$ where both occur in $p$’s resolvent, and a spawned goal $\texttt{global\_send}(X_{p}?,\_w(p,i),q)$
  • At agent $q$: a local variable pair $(X_{q},X_{q}?)$ where both occur in $q$’s resolvent, and an entry $(X_{q},p,i)$ in $q$’s global writers table
Similarly, a global link for a reader (callback) has:
  • At agent $p$: a local pair $(Y_{p},Y_{p}?)$, and an entry $(Y_{p},q)$ at index $i$ in $p$’s global writers table
  • At agent $q$: a local pair $(Z_{q},Z_{q}?)$, and a spawned goal $\texttt{global\_send}(Z_{q}?,\_r(p,i),p)$
Definition C.27 (Variable Correspondence).
A global link from $p$ to $q$ as in Definition C.26 implements the maGLP shared variable pair $(X,X?)$ where writer $X$ is held by $p$ and reader $X?$ is held by $q$. The correspondence is:
  • The madGLP writer $X_{p}$ corresponds to the maGLP writer $X$ at $p$
  • The madGLP reader $X_{q}?$ corresponds to the maGLP reader $X?$ at $q$
  • The readers $X_{p}?$ and writers $X_{q}$ are implementation artifacts that enable communication
C.9 Correctness Proofs
Lemma C.28 (Globalize-Localize Correspondence).
Let $T$ be a term with variables at agent $p$, let $T_{p}^{\uparrow}$ be its globalization by $p$ for remote agent $q$, and let $T_{q}^{\downarrow}$ be the localization by $q$. Then:
  1. 1.
    For each writer $Y$ in $T$: globalization spawns $\texttt{global\_send}(Y?,\_w(p,i),q)$ at $p$; localization creates fresh pair $(Y_{q},Y_{q}?)$ at $q$ with entry $(Y_{q},p,i)$. These form a global link implementing shared pair $(Y,Y?)$.
  2. 2.
    For each reader $Y?$ in $T$: globalization creates entry $(Y,q)$ at index $i$ in $p$’s global writers table; localization creates fresh pair $(Z_{q},Z_{q}?)$ at $q$ and spawns $\texttt{global\_send}(Z_{q}?,\_r(p,i),p)$. These form a global link implementing shared pair $(Y,Y?)$.
  3. 3.
    The structure of $T$ is preserved: $T_{q}^{\downarrow}$ has the same functor/arity structure as $T$, with each variable replaced by its corresponding local variable at $q$.
Proof C.29.
By the definitions of Globalize (Definition C.12) and Localize (Definition C.13):
For (1): When $Y$ is a writer in $T$, globalization replaces $Y$ with _w(p,i)\_w(p,i) and spawns $\texttt{global\_send}(Y?,\_w(p,i),q)$. Localization at $q$ sees _w(p,i)\_w(p,i), creates fresh pair $(Y_{q},Y_{q}?)$, creates entry $(Y_{q},p,i)$, and replaces _w(p,i)\_w(p,i) with $Y_{q}?$ in $T_{q}^{\downarrow}$. When $Y$ is assigned at $p$, the value becomes available via $Y?$, the global_send goal fires and sends message $w(p,i):=T^{\uparrow}$ to $q$. Agent $q$’s Receive searches for entry $(Y_{q},p,i)$ and assigns $Y_{q}:=T^{\downarrow}$, making the value available via $Y_{q}?$. This matches maGLP semantics.
For (2): When $Y?$ is a reader in $T$, globalization replaces $Y?$ with _r(p,i)\_r(p,i) and creates entry $(Y,q)$ at index $i$ in $p$’s global writers table. Localization at $q$ creates fresh pair $(Z_{q},Z_{q}?)$, spawns $\texttt{global\_send}(Z_{q}?,\_r(p,i),p)$, and replaces _r(p,i)\_r(p,i) with $Z_{q}$ in $T_{q}^{\downarrow}$. When $q$ assigns $Z_{q}:=T$, the global_send goal fires and sends $r(p,i):=T^{\uparrow}$ to $p$. Agent $p$’s Receive finds entry $(Y,q)$ at index $i$ and assigns $Y:=T^{\downarrow}$, making the value available via $Y?$. This implements callback semantics.
For (3): Both globalization and localization traverse $T$ structurally, replacing only variables while preserving all functors and their arities.
Lemma C.30 (Both Ends Exported).
When agent $p$ globalizes term $T$ containing both $X$ and $X?$ for agent $q$, the resulting madGLP configuration correctly implements the maGLP semantics where both ends of the pair are available to $q$.
Proof C.31.
Globalization of $T=[\ldots,X,\ldots,X?,\ldots]$ at $p$ for $q$ proceeds as follows:
  • For writer $X$: spawns $\texttt{global\_send}(X?,\_w(p,1),q)$, no entry, term gets _w(p,1)\_w(p,1)
  • For reader $X?$: entry $(X,q)$ at index 2, no spawn, term gets _r(p,2)\_r(p,2)
Localization at $q$ creates:
  • For _w(p,1)\_w(p,1): pair $(Y_{q},Y_{q}?)$, entry $(Y_{q},p,1)$, term gets $Y_{q}?$
  • For _r(p,2)\_r(p,2): pair $(Z_{q},Z_{q}?)$, spawns $\texttt{global\_send}(Z_{q}?,\_r(p,2),p)$, term gets $Z_{q}$
The term at $q$ is $\ldots,Y_{q}?,\ldots,Z_{q},\ldots]$ with two independent pairs. Agent $q$ does not know they are connected—that connection exists only through $p$’s local pair $(X,X?)$.
When $q$ assigns $Z_{q}:=T$:
  1. 1.
    $Z_{q}?$ becomes known
  2. 2.
    $\texttt{global\_send}(Z_{q}?,\_r(p,2),p)$ fires, sends $r(p,2):=T^{\uparrow}$ to $p$
  3. 3.
    $p$’s Receive finds entry $(X,q)$ at index 2 and assigns $X:=T^{\downarrow}$
  4. 4.
    $X?$ becomes known at $p$
  5. 5.
    $\texttt{global\_send}(X?,\_w(p,1),q)$ fires, sends $w(p,1):=T^{\uparrow}$ to $q$
  6. 6.
    $q$’s Receive searches for entry $(Y_{q},p,1)$ and assigns $Y_{q}:=T^{\downarrow}$
  7. 7.
    $Y_{q}?$ becomes known at $q$
The assignment to $Z_{q}$ flows through $p$’s local pair to $Y_{q}?$. The same analysis applies when $X$ and $X?$ are sent to different agents.
The Implementation Mapping.
Definition C.32 (Reconstructed Resolvent).
Given a madGLP local state $(R_{p},W_{p},M_{p})$ where $R_{p}=(A_{p},S_{p},F_{p})$, the reconstructed resolvent at agent $p$ is:
$G_{p}=A_{p}\cup\{A:(A,W)\in S_{p}\}\cup F_{p}$
Definition C.33 (Reconstructed Readers Substitution).
Given a madGLP configuration $c^{\prime}$, the reconstructed readers substitution $\sigma_{p}$ for agent $p$ is derived from in-flight messages: for each message $m=(G:=T^{\uparrow},p)$ in any communication channel destined for $p$, include the corresponding reader assignment in $\sigma_{p}$ according to the variable correspondence.
Definition C.34 (Implementation Mapping $\pi$).
The implementation mapping $\pi$ from madGLP configurations to maGLP configurations is defined as follows. For each agent $p$:
$\pi_{p}(R_{p},W_{p},M_{p})=(G_{p},\sigma_{p})$
where $G_{p}$ is the reconstructed resolvent (Definition C.32) with global_send goals removed and each local variable replaced by its corresponding maGLP variable according to the variable correspondence (Definition C.27), and $\sigma_{p}$ is the reconstructed readers substitution (Definition C.33).
The global mapping $\pi:C^{\prime}\rightarrow C$ applies this construction to each agent’s local state.
The mapping $\pi$ forgets:
  • The partition of goals into active, suspended, and failed
  • The suspension sets $W$ for suspended goals
  • The global writers tables $W_{p}$ (routing information)
  • The message queues $M_{p}$ (except as they contribute to $\sigma$)
  • The global_send goals (implementation mechanism)
  • The pairing of local variables—corresponding variables at different agents map to the same maGLP variable
Transaction Correspondence.
Proposition C.35 (Reduce Implementation).
The madGLP Reduce transaction for agent $p$ projects via $\pi$ to a maGLP Reduce transaction for $p$.
Proof C.36.
[…]
Proposition C.37 (Communicate Implementation).
maGLP’s binary Communicate transaction from $p$ to $q$ is correctly implemented by madGLP’s mechanism: a global_send goal fires (via Reduce), then Send, then Receive at $q$.
Proof C.38.
maGLP Communicate(p,q)(p,q) transfers assignment $\{X?:=T\}$ from $\sigma_{p}$ to $G_{q}$. In madGLP: (1) $p$’s Reduce assigns $X:=T$; (2) a subsequent Reduce fires global_send(X?,G,q)(X?,G,q), adding message $(G:=T^{\uparrow},q)$ to $M_{p}$; (3) Send moves the message to the channel; (4) Receive at $q$ localizes and applies the assignment. By Lemma C.28, $\pi$ maps the result to the maGLP Communicate result.
Proposition C.39 (Cold-call Implementation).
maGLP’s binary Cold-call transaction from $p$ to $q$ is correctly implemented by madGLP’s Send and Receive transactions via the index-0 serializer.
Proof C.40.
maGLP Cold-call(p,q)(p,q) sends term $T$ establishing shared variable pairs. In madGLP: $p$ globalizes $T$ to $T_{p}^{\uparrow}$; Send at $p$ places the message in the channel; Receive at $q$ localizes to $T_{q}^{\downarrow}$; global links are established. By Lemma C.28, $\pi$ maps the result to the maGLP Cold-call result.
Lemma C.41 ($\pi$ is Live).
The implementation mapping $\pi$ is live.
Proof C.42.
By Remark B.3, maGLP is persistent. Each agent’s madGLP execution is deterministic at the local level (FIFO scheduling). For any fair madGLP run $r^{\prime}$, every madGLP Reduce that becomes enabled is eventually taken. Each such Reduce projects to a maGLP Reduce via $\pi$. Since madGLP can perform any set of reductions that maGLP can (by Lemma C.43), and FIFO scheduling ensures all enabled reductions are eventually taken, $\pi(r^{\prime})$ is a fair maGLP run.
Completeness.
Lemma C.43 (madGLP Performs Same Reductions).
For every proper maGLP run $\rho$ performing reductions $\{R_{1},\ldots,R_{k}\}$, there exists a madGLP run $\rho^{\prime}$ performing the same reductions.
Proof C.44.
By induction on spawn order. Goals in each agent’s initial goal are in the initial queue. For any $R_{i}$ reducing $A_{i}$ at agent $p$ with $C_{i}$: if $A_{i}$ was spawned by an earlier reduction $R_{j}$, by induction madGLP performs $R_{j}$, so $A_{i}$ enters $p$’s queue. Since $R_{i}$ succeeds in $\rho$, $A_{i}$ and the head of $C_{i}$ have a writer mgu. By Persistence (Lemma 3.12), this mgu exists regardless of readers substitution. When $A_{i}$ reaches the queue head, the reduction succeeds.
Lemma C.45 (madGLP Completeness).
For every proper maGLP run with outcome $O$, there exists a madGLP run with outcome $O$.
Proof C.46.
Let $\rho$ be a proper maGLP run with outcome $O$, performing reductions $\{R_{1},\ldots,R_{k}\}$.
By Lemma C.43, some madGLP run $\rho^{\prime}$ performs the same reductions.
By Lemma 3.20 and Lemma 3.21, since $\rho$ and $\rho^{\prime}$ perform the same reductions (with disjoint substitutions that commute), they have the same outcome $O$.
Appendix D madGLP is Grassroots
This section establishes that madGLP is grassroots. We first define protocols and the grassroots property, then prove that madGLP satisfies it. We conclude with a general theorem showing that correct and complete implementations preserve the grassroots property, yielding as a corollary that maGLP is grassroots.
D.1 Protocols and the Grassroots Property
A protocol is a family of multiagent transition systems, one for each finite set of agents $P\subset\Pi$, sharing an underlying set of local states with a designated initial state.
Definition D.1 (Local-States Function).
A local-states function $S:2^{\Pi}\to 2^{\mathcal{S}}$ maps every finite set of agents $P\subset\Pi$ to a set of local states $S(P)\subseteq\mathcal{S}$ that includes $s0$ and satisfies $P\subset P^{\prime}\implies S(P)\subseteq S(P^{\prime})$.
Given a local-states function $S$, we use $C(P):=S(P)^{P}$ for configurations over $P$ and $c0(P):=\{s0\}^{P}$ for the initial configuration.
Definition D.2 (Protocol).
A protocol $\mathcal{F}$ over a local-states function $S$ is a family of multiagent transition systems with exactly one transition system $\mathcal{F}(P)=(C(P),c0(P),T(P))$ for every finite $P\subset\Pi$.
Definition D.3 (Projection).
Let $\emptyset\subset P\subset P^{\prime}\subset\Pi$. If $c^{\prime}$ is a configuration over $P^{\prime}$, then $c^{\prime}/P$, the projection of $c^{\prime}$ onto $P$, is the configuration $c$ over $P$ defined by $c_{p}:=c^{\prime}_{p}$ for every $p\in P$.
Note that in the projection $c^{\prime}/P$, each local state $c_{p}$ is in $S(P^{\prime})$, not necessarily in $S(P)$, and hence may include elements “alien” to $P$, such as variables shared with agents in $P^{\prime}\setminus P$.
Definition D.4 (Oblivious, Interactive, Grassroots).
A protocol $\mathcal{F}$ is:
  1. 1.
    oblivious if for every $\emptyset\subset P\subset P^{\prime}\subset\Pi$, $T(P){\uparrow}P^{\prime}\subseteq T(P^{\prime})$
  2. 2.
    interactive if for every $\emptyset\subset P\subset P^{\prime}\subset\Pi$ and every configuration $c\in C(P^{\prime})$ such that $c/P\in C(P)$, there is a computation $c\xrightarrow{*}c^{\prime}$ of $\mathcal{F}(P^{\prime})$ for which $c^{\prime}/P\notin C(P)$
  3. 3.
    grassroots if it is oblivious and interactive
Being oblivious means that agents in $P$ can always perform locally any transition they could perform in isolation, regardless of the presence of additional agents in $P^{\prime}\setminus P$. Being interactive means that agents in $P$ can always eventually interact with agents outside $P$ in a way that leaves “alien traces” in their local states.
D.2 Transactions-Based Protocols
Definition D.5 (Transactions Over a Local-States Function).
Let $S$ be a local-states function. A set of transactions $R$ is over $S$ if every transaction $t\in R$ is a transition over participants $Q$ and local states $S(Q^{\prime})$ for some $Q\subseteq Q^{\prime}\subset\Pi$. Given such $R$ and $P\subset\Pi$:
[…]
Definition D.6 (Transactions-Based Protocol).
Let $S$ be a local-states function and $R$ a set of transactions over $S$. The protocol $\mathcal{F}$ over $R$ and $S$ is defined by $\mathcal{F}(P):=(C(P),c0(P),R(P){\uparrow}P)$ for each $P\subset\Pi$.
Proposition D.7 (Transactions-Based Protocols are Oblivious).
A transactions-based protocol is oblivious.
Proof D.8.
Let $\mathcal{F}$ be a protocol over transactions $R$ and local-states function $S$. Let $P\subset P^{\prime}\subset\Pi$. We must show $T(P){\uparrow}P^{\prime}\subseteq T(P^{\prime})$.
By Definition D.6, $T(P)=R(P){\uparrow}P$. A transition $t^{\prime}\in T(P){\uparrow}P^{\prime}$ is derived from some $t\in R(P)$ by extending to agents in $P^{\prime}$ who remain stationary. Since $t\in R(P)$ has participants $Q\subseteq P\subset P^{\prime}$ and is over $S(Q^{\prime})$ for some $Q^{\prime}\subseteq P\subset P^{\prime}$, we have $t\in R(P^{\prime})$. Hence $t^{\prime}\in R(P^{\prime}){\uparrow}P^{\prime}=T(P^{\prime})$.
Definition D.9 (Interactive Transactions).
A set of transactions $R$ over a local-states function $S$ is interactive if for every $\emptyset\subset P\subset P^{\prime}\subset\Pi$ and every configuration $c\in C(P^{\prime})$ such that $c/P\in C(P)$, there is a computation $c\xrightarrow{*}c^{\prime}$ in $R(P^{\prime}){\uparrow}P^{\prime}$ for which $c^{\prime}/P\notin C(P)$.
Theorem D.10 (Interactive Transactions Induce Grassroots Protocols).
A protocol over an interactive set of transactions is grassroots.
Proof D.11.
Let $\mathcal{F}$ be a protocol over transactions $R$, where $R$ is interactive. By Proposition D.7, $\mathcal{F}$ is oblivious. By Definition D.9, $\mathcal{F}$ is interactive. By Definition D.4, $\mathcal{F}$ is grassroots.
D.3 madGLP is Grassroots
Theorem D.12 (madGLP is Grassroots).
The madGLP protocol is grassroots.
Proof D.13.
By Theorem D.10, it suffices to show that madGLP is a transactions-based protocol with interactive transactions.
madGLP is transactions-based: By Definition 5.6, madGLP is defined via transactions (Reduce, Send, Receive) over a local-states function (deterministic resolvents with global writers tables and message queues) with a common initial state. Hence madGLP is a transactions-based protocol.
[…]
Since $c/P\in C(P)$, the local states of agents in $P$ contain no global links to agents in $P^{\prime}\setminus P$—all their variable pairs are internal to $P$.
Consider agents $p\in P$ and $q\in P^{\prime}\setminus P$. Agent $p$ can execute a Reduce transaction that outputs $\mathit{msg}(q,X)$ on its network output stream, where $X$ is a fresh writer. This triggers $p$’s global_send goal for the index-0 serializer, and a subsequent Send transaction at $p$ places the message in the communication channel to $q$. A Receive transaction at $q$ then processes the message, establishing a global link: agent $q$ creates an entry $(X_{q},p,i)$ in its global writers table and adds $X_{q}?$ to its network input stream.
Agent $q$ can then execute Reduce transactions that eventually assign $X_{q}:=T$ for some term $T$. This triggers $q$’s global_send goal, and Send at $q$ followed by Receive at $p$ delivers the value back.
In either case, the resulting configuration $c^{\prime}$ has agent $p$’s global writers table or spawned goals containing references to agent $q\in P^{\prime}\setminus P$. These constitute “alien traces”—the configuration $c^{\prime}/P$ contains global names or entries referencing agents outside $P$, which could not have been produced by agents in $P$ running alone. Hence $c^{\prime}/P\notin C(P)$.
Since such a computation exists for any valid configuration $c$, the madGLP transactions are interactive.
D.4 Implementations Preserve Grassroots
We now establish a general result: correct and complete implementations preserve the grassroots property. This provides an alternative path to proving a specification is grassroots—prove that its implementation is grassroots.
Theorem D.14 (Implementations Preserve Grassroots).
Let $\mathcal{F}$ be a protocol (the specification) and $\mathcal{F}^{\prime}$ a protocol (the implementation) with a family of implementation mappings $\pi_{P}:C^{\prime}(P)\to C(P)$ such that $(\mathcal{F}^{\prime}(P),\pi_{P})$ is a correct and complete implementation of $\mathcal{F}(P)$ for every $P\subset\Pi$. If $\mathcal{F}^{\prime}$ is grassroots, then $\mathcal{F}$ is grassroots.
Proof D.15.
We show that $\mathcal{F}$ is oblivious and interactive.
Oblivious: Let $P\subset P^{\prime}\subset\Pi$ and let $t\in T(P)$ be a transition of $\mathcal{F}(P)$. We must show $t{\uparrow}P^{\prime}\subseteq T(P^{\prime})$.
By completeness of $(\mathcal{F}^{\prime}(P),\pi_{P})$, transition $t$ has a corresponding computation in $\mathcal{F}^{\prime}(P)$. Since $\mathcal{F}^{\prime}$ is oblivious, this computation lifts to $\mathcal{F}^{\prime}(P^{\prime})$. By correctness of $(\mathcal{F}^{\prime}(P^{\prime}),\pi_{P^{\prime}})$, the lifted computation projects to a computation in $\mathcal{F}(P^{\prime})$. The projected computation includes $t$ extended to $P^{\prime}$ (agents in $P^{\prime}\setminus P$ remain stationary). Hence $t{\uparrow}P^{\prime}\subseteq T(P^{\prime})$.
Interactive: Let $P\subset P^{\prime}\subset\Pi$ and let $c\in C(P^{\prime})$ with $c/P\in C(P)$. We must show there exists a computation $c\xrightarrow{*}c^{\prime}$ in $\mathcal{F}(P^{\prime})$ with $c^{\prime}/P\notin C(P)$.
By completeness, there exists $c^{*}\in C^{\prime}(P^{\prime})$ with $\pi_{P^{\prime}}(c^{*})=c$. Since $\mathcal{F}^{\prime}$ is interactive, there is a computation $c^{*}\xrightarrow{*}c^{*\prime}$ in $\mathcal{F}^{\prime}(P^{\prime})$ such that $c^{*\prime}/P\notin C^{\prime}(P)$.
By correctness, $\pi_{P^{\prime}}(c^{*}\xrightarrow{*}c^{*\prime})$ is a computation $c\xrightarrow{*}c^{\prime}$ in $\mathcal{F}(P^{\prime})$. We claim $c^{\prime}/P\notin C(P)$.
Suppose for contradiction that $c^{\prime}/P\in C(P)$. Since $\pi_{P}$ is surjective (by completeness), there exists $d\in C^{\prime}(P)$ with $\pi_{P}(d)=c^{\prime}/P$. But $\pi_{P^{\prime}}(c^{*\prime})/P=c^{\prime}/P=\pi_{P}(d)$, and the implementation mapping respects projections: $\pi_{P^{\prime}}(c^{*\prime})/P=\pi_{P}(c^{*\prime}/P)$. This would imply $c^{*\prime}/P$ maps to something in $C(P)$, suggesting $c^{*\prime}/P\in C^{\prime}(P)$ (since $\pi_{P}$ maps $C^{\prime}(P)$ to $C(P)$). But this contradicts $c^{*\prime}/P\notin C^{\prime}(P)$.
Hence $c^{\prime}/P\notin C(P)$, establishing interactivity.
Corollary D.16 (maGLP is Grassroots).
The maGLP protocol is grassroots.
Proof D.17.
By Theorem 5.7, madGLP correctly and completely implements maGLP, with implementation mapping $\pi$ that respects the protocol structure (each $\pi_{P}$ maps madGLP(P)(P) configurations to maGLP(P)(P) configurations). By Theorem D.12, madGLP is grassroots. By Theorem D.14, maGLP is grassroots.
Remark D.18 (Two Paths to Grassroots).
This section demonstrates two approaches to proving a system is grassroots: (1) directly, by showing it is transactions-based with interactive transactions (Theorem D.12), or (2) indirectly, by showing it is correctly implemented by a grassroots system (Corollary D.16). The indirect approach is useful when the implementation has simpler structure than the specification.
Appendix E Detailed madGLP Example Traces
This appendix provides complete formal traces for three madGLP examples: the client-monitor example (Program 5.2), the friend-mediated introduction scenario (Bob sends $X$ to Alice and $X?$ to Charlie), and the edge case where both ends of a variable pair are sent to the same agent.
E.1 Example 1: Client-Monitor Communication
This example corresponds to Figure 1.
Initial goal: client1(Xs)@p, monitor(Xs?)@q
Stage 0: After Initial Cold-call.
maGLP State
Agent $p$ Resolvent Agent $q$ Resolvent
$\mathsf{Xs}$ $\mathsf{Xs?}$
Shared Pairs: $(\mathsf{Xs}@p,\mathsf{Xs?}@q)$
madGLP State
Agent $p$ Agent $q$
Resolvents (including global_send goals)
$\mathsf{Xs_{p}},\mathsf{Xs_{p}?}$, $\mathsf{Xs_{q}},\mathsf{Xs_{q}?}$
[…]
Bindings
(none) (none)
Global Writers Tables
$k$ entry
0 $(\mathsf{Xs_{q}},p,0)$
Pending Messages: (none)
[…]
Stage 1a: After $p$ Reduces (before $q$ Receives).
Agent $p$ assigns $\mathsf{Xs_{p}}:=[\mathtt{add}|\mathsf{Xs1_{p}}]$. This makes $\mathsf{Xs_{p}?}$ known, triggering the global_send goal.
madGLP State
Agent $p$ Agent $q$
Resolvents (including global_send goals)
$\mathsf{Xs1_{p}},\mathsf{Xs1_{p}?}$, $\mathsf{Xs_{q}},\mathsf{Xs_{q}?}$
[…]
Bindings
$\mathsf{Xs_{p}}\mapsto[\mathtt{add}|\mathsf{Xs1_{p}}]$ (none)
$\mathsf{Xs_{p}?}\mapsto[\mathtt{add}|\mathsf{Xs1_{p}?}]$
Global Writers Tables
$k$ entry
0 $(\mathsf{Xs_{q}},p,0)$
Pending Messages
$(\_w(p,0):=[\mathtt{add}|\_w(p,1)],\ q)$
Globalization performed:
Term before $\mathtt{add}|\mathsf{Xs1_{p}}]$
Term after $\mathtt{add}|\_w(p,1)]$
[…]
New entry at $p$ (none—outgoing link)
The old global_send goal for $\mathsf{Xs_{p}?}$ fired (since $\mathsf{Xs_{p}?}$ became known) and completed after adding the message to $M_{p}$. A new global_send goal was spawned for $\mathsf{Xs1_{p}}$.
Stage 1b: After $q$ Receives.
Agent $q$ receives message and localizes.
maGLP State
Agent $p$ Resolvent Agent $q$ Resolvent
$\mathsf{Xs1}$ $\mathsf{Xs1?}$
Bindings
$\mathsf{Xs}\mapsto[\mathtt{add}|\mathsf{Xs1}]$ $\mathsf{Xs?}\mapsto[\mathtt{add}|\mathsf{Xs1?}]$
Shared Pairs: $(\mathsf{Xs1}@p,\mathsf{Xs1?}@q)$
madGLP State
Agent $p$ Agent $q$
Resolvents (including global_send goals)
$\mathsf{Xs1_{p}},\mathsf{Xs1_{p}?}$, $\mathsf{Xs1_{q}},\mathsf{Xs1_{q}?}$
[…]
Bindings
$\mathsf{Xs_{p}}\mapsto[\mathtt{add}|\mathsf{Xs1_{p}}]$ $\mathsf{Xs_{q}}\mapsto[\mathtt{add}|\mathsf{Xs1_{q}?}]$
$\mathsf{Xs_{p}?}\mapsto[\mathtt{add}|\mathsf{Xs1_{p}?}]$ $\mathsf{Xs_{q}?}\mapsto[\mathtt{add}|\mathsf{Xs1_{q}?}]$
Global Writers Tables
$k$ entry
1 $(\mathsf{Xs1_{q}},p,1)$
Pending Messages: (none)
Localization performed:
Term before $\mathtt{add}|\_w(p,1)]$
Term after $\mathtt{add}|\mathsf{Xs1_{q}?}]$
New pair at $q$ $(\mathsf{Xs1_{q}},\mathsf{Xs1_{q}?})$
New entry at $q$ $(1,(\mathsf{Xs1_{q}},p,1))$
New goal at $q$ (none—incoming link)
Entry 0 at $q$ was removed after assigning $\mathsf{Xs_{q}}$.
Stage 2a: After $p$ Reduces (before $q$ Receives).
Agent $p$ assigns $\mathsf{Xs1_{p}}:=[\mathtt{value}(\mathsf{V_{p}?})|\mathsf{Xs2_{p}}]$ and globalizes.
madGLP State
Agent $p$ Agent $q$
Resolvents (including global_send goals)
$\mathsf{V_{p}},\mathsf{V_{p}?},\mathsf{Xs2_{p}},\mathsf{Xs2_{p}?}$, $\mathsf{Xs1_{q}},\mathsf{Xs1_{q}?}$
[…]
Bindings
$\mathsf{Xs_{p}}\mapsto[\mathtt{add}|\mathsf{Xs1_{p}}]$ $\mathsf{Xs_{q}}\mapsto[\mathtt{add}|\mathsf{Xs1_{q}?}]$
$\mathsf{Xs_{p}?}\mapsto[\mathtt{add}|\mathsf{Xs1_{p}?}]$ $\mathsf{Xs_{q}?}\mapsto[\mathtt{add}|\mathsf{Xs1_{q}?}]$
$\mathsf{Xs1_{p}}\mapsto[\mathtt{value}(\mathsf{V_{p}?})|\mathsf{Xs2_{p}}]$
$\mathsf{Xs1_{p}?}\mapsto[\mathtt{value}(\mathsf{V_{p}?})|\mathsf{Xs2_{p}?}]$
Global Writers Tables
$i$ entry
2 $(\mathsf{V_{p}},q)$
$k$ entry
1 $(\mathsf{Xs1_{q}},p,1)$
Pending Messages
$(\_w(p,1):=[\mathtt{value}(\_r(p,2))|\_w(p,3)],\ q)$
Globalization performed:
Term before $\mathtt{value}(\mathsf{V_{p}?})|\mathsf{Xs2_{p}}]$
Term after $\mathtt{value}(\_r(p,2))|\_w(p,3)]$
[…]
Stage 2b: After $q$ Receives.
Agent $q$ receives message and localizes.
maGLP State
Agent $p$ Resolvent Agent $q$ Resolvent
$\mathsf{V?},\ \mathsf{Xs2}$ $\mathsf{V},\ \mathsf{Xs2?}$
Bindings
$\mathsf{Xs}\mapsto[\mathtt{add}|\mathsf{Xs1}]$ $\mathsf{Xs?}\mapsto[\mathtt{add}|\mathsf{Xs1?}]$
$\mathsf{Xs1}\mapsto[\mathtt{value}(\mathsf{V?})|\mathsf{Xs2}]$ $\mathsf{Xs1?}\mapsto[\mathtt{value}(\mathsf{V})|\mathsf{Xs2?}]$
Shared Pairs: $(\mathsf{V}@q,\mathsf{V?}@p)$, $(\mathsf{Xs2}@p,\mathsf{Xs2?}@q)$
madGLP State
Agent $p$ Agent $q$
Resolvents (including global_send goals)
$\mathsf{V_{p}},\mathsf{V_{p}?},\mathsf{Xs2_{p}},\mathsf{Xs2_{p}?}$, $\mathsf{V_{q}},\mathsf{V_{q}?},\mathsf{Xs2_{q}},\mathsf{Xs2_{q}?}$,
$(\mathsf{Xs2_{p}?},\_w(p,3),q) global_send$$(\mathsf{V_{q}?},\_r(p,2),p)$
Bindings
$\mathsf{Xs_{p}}\mapsto[\mathtt{add}|\mathsf{Xs1_{p}}]$ $\mathsf{Xs_{q}}\mapsto[\mathtt{add}|\mathsf{Xs1_{q}?}]$
$\mathsf{Xs_{p}?}\mapsto[\mathtt{add}|\mathsf{Xs1_{p}?}]$ $\mathsf{Xs_{q}?}\mapsto[\mathtt{add}|\mathsf{Xs1_{q}?}]$
$\mathsf{Xs1_{p}}\mapsto[\mathtt{value}(\mathsf{V_{p}?})|\mathsf{Xs2_{p}}]$ $\mathsf{Xs1_{q}}\mapsto[\mathtt{value}(\mathsf{V_{q}})|\mathsf{Xs2_{q}?}]$
$\mathsf{Xs1_{p}?}\mapsto[\mathtt{value}(\mathsf{V_{p}?})|\mathsf{Xs2_{p}?}]$ $\mathsf{Xs1_{q}?}\mapsto[\mathtt{value}(\mathsf{V_{q}})|\mathsf{Xs2_{q}?}]$
Global Writers Tables
$i$ entry
2 $(\mathsf{V_{p}},q)$
$k$ entry
2 $(\mathsf{Xs2_{q}},p,3)$
Pending Messages: (none)
Localization performed:
Term before $\mathtt{value}(\_r(p,2))|\_w(p,3)]$
Term after $\mathtt{value}(\mathsf{V_{q}})|\mathsf{Xs2_{q}?}]$
[…]
Global Link Summary.
After Stage 2b, the active global links are:
Link Entry at Entry Direction global_send at
V $p$ (index 2) $(\mathsf{V_{p}},q)$ $\to p qq$
Xs2 $q$ (index 2) $(\mathsf{Xs2_{q}},p,3)$ $\to q pp$
The “Direction” column indicates dataflow: V flows from $q$ to $p$, while Xs2 flows from $p$ to $q$ (normal case). Entries exist only at the receiving end; global_send goals exist only at the sending end.
E.2 Example 2: Friend-Mediated Introduction
This example demonstrates the critical case where Bob (agent $b$) introduces Alice (agent $a$) to Charlie (agent $c$) by sending them opposite ends of a variable pair. Bob sends writer $X$ to Alice and reader $X?$ to Charlie. Bob’s local pair $(X_{b},X_{b}?)$ acts as a relay: when Charlie assigns his writer, the value flows through Bob to Alice.
Stage 0: Initial State at Bob.
Bob has a local variable pair $(X_{b},X_{b}?)$ that will be shared with Alice and Charlie.
madGLP State
Agent $a$ (Alice) Agent $b$ (Bob) Agent $c$ (Charlie)
Resolvents
(empty) $X_{b},X_{b}?$ (empty)
Global Writers Tables
(empty) (empty) (empty)
Stage 1: Bob Sends $X$ to Alice via Cold-call.
Bob globalizes $X_{b}$ (a writer) for Alice.
Globalization at Bob for Alice:
Variable $X_{b}$ (writer)
Global name _w(b,0)\_w(b,0)
Entry at Bob (none—outgoing link)
[…]
Localization at Alice:
Global name _w(b,0)\_w(b,0)
New pair $(X_{a},X_{a}?)$
In term $X_{a}?$ (reader)
Entry at Alice $(X_{a},b,0)$
Spawned goal (none—incoming link)
madGLP State after Stage 1
Agent $a$ (Alice) Agent $b$ (Bob) Agent $c$ (Charlie)
Resolvents (including global_send goals)
$X_{a},X_{a}?$ $X_{b},X_{b}?$, (empty)
[…]
Global Writers Tables
$k$ entry
0 $(X_{a},b,0)$
(empty) (empty)
Stage 2: Bob Sends $X?$ to Charlie via Cold-call.
Bob globalizes $X_{b}?$ (a reader) for Charlie.
Globalization at Bob for Charlie:
Variable $X_{b}?$ (reader)
Global name _r(b,1)\_r(b,1)
Entry at Bob $(X_{b},c)$ at index 1
Spawned goal (none—incoming link)
Localization at Charlie:
Global name _r(b,1)\_r(b,1)
New pair $(X_{c},X_{c}?)$
In term $X_{c}$ (writer)
Entry at Charlie (none—outgoing link)
[…]
madGLP State after Stage 2
Agent $a$ (Alice) Agent $b$ (Bob) Agent $c$ (Charlie)
Resolvents (including global_send goals)
$X_{a},X_{a}?$ $X_{b},X_{b}?$, $X_{c},X_{c}?$,
$(X_{b}?,\_w(b,0),a) global_send$$(X_{c}?,\_r(b,1),b)$
Global Writers Tables
$k$ entry
0 $(X_{a},b,0)$
$i$ entry
1 $(X_{b},c)$
Global Link Summary after Stage 2:
Link Entry at Entry global_send at
_w(b,0)\_w(b,0) Alice $(X_{a},b,0)$ Bob
_r(b,1)\_r(b,1) Bob (index 1) $(X_{b},c)$ Charlie
Stage 3: Charlie Assigns $X_{c}:=T$.
Charlie assigns $X_{c}:=T$ for some term $T$. This makes $X_{c}?$ known.
madGLP State after Charlie’s assignment
Agent $a$ (Alice) Agent $b$ (Bob) Agent $c$ (Charlie)
Resolvents
$X_{a},X_{a}?$ $X_{b},X_{b}?$, (goal completed)
[…]
Bindings
(none) (none) $X_{c}\mapsto T$, $X_{c}?\mapsto T$
Pending Messages
$(\_r(b,1):=T^{\uparrow},\ b)$ from Charlie
[…]
Stage 4: Bob Receives from Charlie.
Bob receives message $(\_r(b,1):=T^{\uparrow})$. He finds entry $(X_{b},c)$ at index 1 and assigns $X_{b}:=T^{\downarrow}$.
[…]
madGLP State after Bob receives
Agent $a$ (Alice) Agent $b$ (Bob) Agent $c$ (Charlie)
Resolvents
$X_{a},X_{a}?$ (goal completed) (empty)
Bindings
(none) $X_{b}\mapsto T$, $X_{b}?\mapsto T$ $X_{c}\mapsto T$, $X_{c}?\mapsto T$
Global Writers Tables
$k$ entry
0 $(X_{a},b,0)$
(empty) (empty)
Pending Messages
$(\_w(b,0):=T^{\uparrow},\ a)$ from Bob
Entry at index 1 at Bob was removed after the assignment. The global_send goal fired and sent message $(\_w(b,0):=T^{\uparrow},a)$.
Stage 5: Alice Receives from Bob.
Alice receives message $(\_w(b,0):=T^{\uparrow})$. She searches for entry with remote agent $b$ and remote index 0, finds $(X_{a},b,0)$, and assigns $X_{a}:=T^{\downarrow}$.
Final madGLP State
Agent $a$ (Alice) Agent $b$ (Bob) Agent $c$ (Charlie)
Bindings
$X_{a}\mapsto T$, $X_{a}?\mapsto T$ $X_{b}\mapsto T$, $X_{b}?\mapsto T$ $X_{c}\mapsto T$, $X_{c}?\mapsto T$
Global Writers Tables
(empty) (empty) (empty)
All entries have been removed. The value $T$ has flowed from Charlie through Bob to Alice.
Summary:
Bob acted as a relay. Charlie assigned his writer $X_{c}$, the value flowed to Bob via the _r(b,1)\_r(b,1) link, Bob’s local pair $(X_{b},X_{b}?)$ received it, and the global_send goal watching $X_{b}?$ forwarded it to Alice via the _w(b,0)\_w(b,0) link. Alice’s reader $X_{a}?$ now has the value $T$.
E.3 Example 3: Both Ends to Same Agent
This example demonstrates the edge case where Bob sends both writer $X$ and reader $X?$ to Alice in the same term $[X,X?]$. Bob’s local pair acts as a relay: when Alice assigns her writer (from the reader global name), the value flows through Bob back to Alice’s reader (from the writer global name).
Stage 0: Initial State.
Bob has a local variable pair $(X_{b},X_{b}?)$ and will send term $X_{b},X_{b}?]$ to Alice.
madGLP State
Agent $a$ (Alice) Agent $b$ (Bob)
Resolvents
(empty) $X_{b},X_{b}?$
Global Writers Tables
(empty) (empty)
Stage 1: Bob Sends $X_{b},X_{b}?]$ to Alice via Cold-call.
Bob globalizes term $X_{b},X_{b}?]$ for Alice. Both the writer and reader are processed:
Globalization at Bob:
Variable $X_{b}$ (writer) $X_{b}?$ (reader)
Global name _w(b,0)\_w(b,0) _r(b,1)\_r(b,1)
Entry at Bob (none—outgoing) $(X_{b},a)$ at index 1
[…]
Globalized term:
[_w(b,0),_r(b,1)][\_w(b,0),\_r(b,1)]
Localization at Alice:
Global name _w(b,0)\_w(b,0) _r(b,1)\_r(b,1)
New pair $(Y_{a},Y_{a}?)$ $(Z_{a},Z_{a}?)$
In term $Y_{a}?$ (reader) $Z_{a}$ (writer)
Entry at Alice $(Y_{a},b,0)$ (none—outgoing)
[…]
Localized term:
$Y_{a}?,Z_{a}]$
madGLP State after Stage 1
Agent $a$ (Alice) Agent $b$ (Bob)
Resolvents (including global_send goals)
$Y_{a},Y_{a}?,Z_{a},Z_{a}?$, $X_{b},X_{b}?$,
$(Z_{a}?,\_r(b,1),b) global_send$$(X_{b}?,\_w(b,0),a)$
Global Writers Tables
$k$ entry
0 $(Y_{a},b,0)$
$i$ entry
1 $(X_{b},a)$
Alice’s term is $Y_{a}?,Z_{a}]$: she has reader $Y_{a}?$ (will receive values) and writer $Z_{a}$ (can assign values). These are two independent pairs at Alice—she doesn’t know they’re connected through Bob.
Stage 2: Alice Assigns $Z_{a}:=T$.
Alice assigns $Z_{a}:=T$. This makes $Z_{a}?$ known, triggering her global_send goal.
madGLP State after Alice’s assignment
Agent $a$ (Alice) Agent $b$ (Bob)
Resolvents
$Y_{a},Y_{a}?$ $X_{b},X_{b}?$,
[…]
Bindings
$Z_{a}\mapsto T$, $Z_{a}?\mapsto T$ (none)
Global Writers Tables
$k$ entry
0 $(Y_{a},b,0)$
$i$ entry
1 $(X_{b},a)$
Pending Messages
$(\_r(b,1):=T^{\uparrow},\ b)$ from Alice
The global_send goal sent message $(\_r(b,1):=T^{\uparrow},b)$.
Stage 3: Bob Receives from Alice.
Bob receives message $(\_r(b,1):=T^{\uparrow})$. He finds entry $(X_{b},a)$ at index 1 and assigns $X_{b}:=T^{\downarrow}$.
[…]
madGLP State after Bob receives
Agent $a$ (Alice) Agent $b$ (Bob)
Resolvents
$Y_{a},Y_{a}?$ (goal completed)
Bindings
$Z_{a}\mapsto T$, $Z_{a}?\mapsto T$ $X_{b}\mapsto T$, $X_{b}?\mapsto T$
Global Writers Tables
$k$ entry
0 $(Y_{a},b,0)$
Pending Messages
$(\_w(b,0):=T^{\uparrow},\ a)$ from Bob
Entry at index 1 at Bob was removed after the assignment. The global_send goal fired and sent message $(\_w(b,0):=T^{\uparrow},a)$.
Stage 4: Alice Receives from Bob.
Alice receives message $(\_w(b,0):=T^{\uparrow})$. She searches for entry with remote agent $b$ and remote index 0, finds $(Y_{a},b,0)$, and assigns $Y_{a}:=T^{\downarrow}$.
Final madGLP State
Agent $a$ (Alice) Agent $b$ (Bob)
Bindings
$Y_{a}\mapsto T$, $Y_{a}?\mapsto T$ $X_{b}\mapsto T$, $X_{b}?\mapsto T$
$Z_{a}\mapsto T$, $Z_{a}?\mapsto T$
Global Writers Tables
(empty) (empty)
All entries have been removed.
Summary:
Alice received term $Y_{a}?,Z_{a}]$ with two independent pairs. When she assigned $Z_{a}:=T$, the value flowed to Bob via the _r(b,1)\_r(b,1) link. Bob’s local pair received it, and his global_send goal forwarded it back to Alice via the _w(b,0)\_w(b,0) link. Now both $Y_{a}?$ and $Z_{a}?$ equal TT—the two pairs at Alice are semantically linked through Bob’s relay, even though Alice sees them as independent.
This correctly implements the maGLP semantics: in maGLP, Alice would have received $[X?,X]$ where both variables refer to the same shared pair. Assigning the writer makes the value available to the reader. In madGLP, this is achieved through Bob’s relay.